### v1

# AI Agentic Programming: A Survey of Techniques, Challenges, and Opportunities

Abstract.
AI agentic programming is an emerging paradigm in which large language models (LLMs) autonomously plan, execute, and interact with external tools like compilers, debuggers, and version control systems to iteratively perform complex software development tasks. Unlike conventional code generation tools, agentic systems are capable of decomposing high-level goals, coordinating multi-step processes, and adapting their behavior based on intermediate feedback. These capabilities are transforming the software development practice. As this emerging field evolves rapidly, there is a need to define its scope, consolidate its technical foundations, and identify open research challenges. This survey provides a comprehensive and timely review of AI agentic programming. We introduce a taxonomy of agent behaviors and system architectures, and examine core techniques including planning, memory and context management, tool integration, and execution monitoring. We also analyze existing benchmarks and evaluation methodologies used to assess coding agent performance. Our study identifies several key challenges, including limitations in handling long context, a lack of persistent memory across tasks, and concerns around safety, alignment with user intent, and collaboration with human developers. We discuss emerging opportunities to improve the reliability, adaptability, and transparency of agentic systems. By synthesizing recent advances and outlining future directions, this survey aims to provide a foundation for research and development in building the next generation of intelligent and trustworthy AI coding agents.
Large Language Models, LLMs, AI Agents, AI Agentic Programming
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software development techniques
1. Introduction
The paradigm of software development is changing rapidly with the rise of large language models (LLMs) (hou2024largelanguagemodelssoftware, ). These models enable artificial intelligence (AI) systems that not only generate code (TOSEM25, ) but also understand task requirements, interact with development tools, and iteratively refine their outputs (FSE24_Alshahwan, ; qian2024chatdevcommunicativeagentssoftware, ). Recent studies suggest that software developers now use LLMs routinely to assist in daily coding tasks (hou2024largelanguagemodelssoftware, ; DBLP:conf/fose-ws/FanGHLSYZ23, ; gong2025language, ). Unlike traditional code generation tools (TraditionalCodeGenration, ) that respond to a single prompt with a static code snippet, emerging AI coding agents are designed to operate within dynamic software environments, performing iterative, tool-augmented tasks to achieve complex goals.
This shift has given rise to a new programming paradigm known as AI agentic programming, where LLM-based coding agents can autonomously plan, execute, and refine software development tasks (gao2023retrieval, ; novikov2025alphaevolve, ). These agents go beyond code completion: they can generate entire programs or modules from natural language specifications, diagnose and fix bugs using compiler or test feedback, write and execute test cases, and refactor code for readability or performance. They can also invoke and interact with external tools such as compilers, debuggers, performance profilers, or version control systems, supporting an end-to-end development workflow.
This emerging paradigm has the potential to fundamentally change how software is built and maintained. For example, an AI agent can take a natural language description of a feature and work through a series of steps, such as writing code, generating tests, running those tests, analyzing and fixing issues, and preparing a pull request. Some state-of-the-art coding agents have demonstrated the ability to continue working for hours while maintaining task consistency, avoiding deadlocks, and recovering from failed actions (qian2024chatdevcommunicativeagentssoftware, ; novikov2025alphaevolve, ). These systems can generate and test code, migrate software between frameworks, debug runtime failures, and integrate new features by decomposing complex goals into manageable subtasks (deng2024pentestgptllmempoweredautomaticpenetration, ; liu2023bolaa, ). This represents a clear shift from static, one-shot AI-based code generation to interactive, iterative, and tool-augmented workflows.
Although progress has been fast, AI agentic programming is still in its early stages. Existing systems vary in architecture, autonomy, tool integration, and reasoning capabilities. There is no standard taxonomy, benchmark suite, or evaluation methodology. Moreover, multiple key challenges remain, including improving reliability (hou2024largelanguagemodelssoftware, ), reducing errors or hallucinations (gong2025language, ), handling tasks across different platforms and languages (DBLP:conf/pldi/polycoder22, ), and making sure these systems are safe and trustworthy in real use (yang2024swe, ).
The success of AI coding agents depends heavily on their ability to interact effectively with external tools. However, today’s programming languages, compilers, and debuggers are fundamentally human-centric. They are not designed for automated, autonomous systems. These tools often abstract away internal states and decision-making processes to improve usability, ensure portability, and reduce cognitive load for human users (llvm, ; Maleki, ). While this abstraction benefits human developers, it may not fit AI agents, which require fine-grained, structured access to internal states, transformation sequences, and validation logic in order to reason about the effects of their actions (bi2024iterative, ). Without such access, AI agents struggle to diagnose failures, understand the implications of their changes, or recover from errors in a principled way. For instance, when a code transformation leads to a build failure, the agent needs more than just an error message - it must trace the failure to specific intermediate steps and understand why the changes caused the issue. Existing development environments do not provide hooks and feedback mechanisms to support this kind of iterative, tool-integrated reasoning.
Similarly, agentic coding systems benefit significantly from toolchains that support iterative development, state tracking, and rich feedback propagation - capabilities that most conventional tools do not expose. To operate effectively, AI agents may need access to internal compiler representations, transformation traces, symbolic information, and execution metadata. This raises a fundamental question: Are our current programming languages and software development tools still adequate in the era of AI agentic programming? Or is it time to rethink the design of programming languages, compilers, and debuggers to treat AI agents as first-class participants in the development process?
These challenges show that AI agentic programming is not just a new way of using existing tools. It is a shift that exposes important gaps in how today’s software systems are designed. As the field evolves rapidly, there is a growing need to clarify its conceptual landscape, identify common patterns and system architectures, and assess the suitability of current development ecosystems. This is the right moment to step back, take stock of recent progress, and lay out the key questions that researchers and developers need to tackle next.
Therefore, this survey aims to provide a comprehensive overview of the emerging field of AI agentic programming. Specifically, it covers:
  • A conceptual foundation and taxonomy of AI coding agents,
  • A review of core system architectures and underlying techniques,
  • A summary of current applications and practical use cases,
  • An analysis of evaluation strategies and benchmarking practices,
  • A discussion of key challenges and current limitations, and
  • An exploration of future research directions, including opportunities to bridge perspectives across disciplines such as programming languages, software engineering, AI, and human-computer interaction.
We focus primarily on LLM-driven agentic systems for software development, though many insights extend to general task-oriented agents. Our goal is to chart the current landscape, clarify foundational concepts, and support the design of robust, efficient, and trustworthy AI agents for programming.
2. Background
2.1. AI Agentic Programming
Figure 1. A representative workflow of an AI coding agent.
AI agentic programming refers to a new programming paradigm in which LLM-based agents autonomously perform software development tasks. Unlike traditional code generation tools that produce outputs in a single step based on a static prompt (DBLP:conf/pldi/polycoder22, ), agentic systems operate in a goal-directed, multi-step manner. They reason about tasks, make decisions, use external tools (such as compilers, debuggers, and test runners), and iteratively refine their outputs based on feedback (DBLP:conf/nips/ShinnCGNY23, ; liu2024large, ; novikov2025alphaevolve, ). These agents can plan sequences of actions, adapt their strategies over time, and coordinate complex development workflows with limited or no human intervention.
At its core, agentic programming combines the capabilities from natural language processing, external tool integration, and task planning. Figure 1 illustrates the architecture of a Copilot-style agentic programming system. At its core, the agent embeds an LLM within an execution loop, enabling interaction with the development environment. The agent receives natural language prompts from the user and gathers additional context from the operating system and the workspace (e.g., file summaries or environment state). This information is passed into the reasoning loop, where the LLM decomposes the task into subgoals, generates code or decisions, and determines whether to invoke external tools—such as reading/editing files or executing terminal commands. Tool outputs are returned to the loop and used as feedback for further refinement. This iterative process continues until the agent completes the task or reaches a stopping condition. Final results are streamed back to the user.
AI agentic programming is characterized by several key properties. First, it emphasizes autonomy, where LLM-based agents can make decisions and take actions without continuous human supervision. Second, it is inherently interactive, as agents engage with external tools and environments during execution. Third, it supports iterative refinement, allowing agents to improve outputs based on intermediate feedback. Finally, it is goal-oriented, with agents pursuing high-level objectives (e.g., sub-tasks generated from the user inputs) rather than simply responding to one-shot prompts.
Together, these features mark a departure from earlier forms of automation and code generation based on rules (xu2008rule, ), classical machine learning models (TraditionalCodeGenration, ) or one-shot LLM calling (DBLP:conf/pldi/polycoder22, ). AI agentic programming represents a change toward intelligent systems that actively participate in the software development process. This enables new capabilities in intelligent code assistance (github_copilot, ), autonomous debugging and testing (englhardt2024exploring, ; deng2024pentestgptllmempoweredautomaticpenetration, ), automated code maintenance (fan2023automated, ), and potentially even self-improving software systems (DBLP:journals/corr/abs-2405-15189, ; DBLP:journals/corr/abs-2310-02304, ).
2.1.1. Working example
Figure 2. Agentic workflow for implementing a REST task.
As an example of AI agentic programming, consider a developer who tasks an AI coding agent with the following request: “Implement a REST API endpoint that returns the top 10 most frequently accessed URLs from a web server log file. Include unit tests and documentation.” This task requires integrating multiple software components, including file parsing, frequency analysis, web API implementation, testing, and documentation.
A high-level view of an agentic loop for solving this task is depicted in Figure 2. Here, an LLM begins by analyzing the natural language task and planning a sequence of actions. It first produces a Python function to parse the log file and count URL frequencies using a dictionary or a data analysis library like collections.Counter. Next, it implements a REST API endpoint using a web framework such as Flask, exposing a route like /top-urls that returns the computed result in JSON format. The LLM then writes unit tests and calls a Python interpreter to execute the generated Python script in the terminal and collects output to validate both the parsing logic and the API usage. It runs these tests using a tool like pytest, identifies failing cases, and refines the implementation. If a test fails due to a corner case (e.g., missing fields or malformed input), the LLM goes back to change the Python code, e.g., by adding input validation. It repeats this process - running tools, interpreting results, and modifying the code - until all tests pass. Finally, the LLM can generate documentation strings for each function and call an external tool like pdoc or Sphinx to produce human-readable API documentation. The process concludes when the agent validates that the tests pass, the API behaves as expected, and the documentation is complete.
This example illustrates the core features of agentic programming: autonomous planning, tool integration, iterative refinement, and goal-directed behavior. Unlike one-shot code generation, the agent interacts continuously with tools, learns from feedback, and adapts its actions to deliver a complete and functional software component.
2.2. Historical Context and Motivations
Figure 3. A diagram showing the evolution of coding agents from program synthesis, to code completion tools to AI coding agents
As illustrated in Figure 3, the idea of automating software development has long been a goal in artificial intelligence and software engineering research (solar2008program, ). Early efforts in program synthesis aimed to generate correct-by-construction programs from formal specifications (gulwani2017program, ), while code completion tools sought to improve developer productivity by predicting likely code snippets based on contexts (github_copilot, ; tabnine, ; amazon-q-developer, ). These approaches, while impactful, typically relied on classical machine-learning models (TraditionalCodeGenration, ), handcrafted ruless (xu2008rule, ), or statistical techniques with limited generalizations (bruch2009learning, ).
The advent of large-scale pre-trained LLMs such as Codexs (openai2021codex, ), GPTs (openai2024gpt4o, ), Claudes (claude2024, ), LLaMas (meta2024llama3, ) and StarCoder (lozhkov2024starcoder, ) marked a major turning point. These models demonstrated strong zero-shot and few-shot capabilities in generating code (DBLP:conf/nips/Le0GSH22, ; peng2024perfcodegen, ), translating between programming languages (DBLP:conf/emnlp/2020pymt5, ; DBLP:conf/mipro/GuzziKPGB24, ), and answering complex programming questions with little or no task-specific fine-tuning (DBLP:journals/corr/abs-2203-07814, ; coignion2024performance, ). Their ability to understand and generate natural language and code made them a good fit for software development tasks beyond basic code completion (husein2025large, ), including documentation generation (dvivedi2024comparative, ), test synthesis (jorgensen2024large, ), and bug detection and fixing (englhardt2024exploring, ; deng2024pentestgptllmempoweredautomaticpenetration, ; kang2025explainable, ).
As these models became more capable, a new opportunity emerged: using LLMs not just as passive code generators, but as autonomous agents that could reason about goals, invoke tools, and refine their outputs over multiple steps. This shift leads to the paradigm of AI agentic programming, where models operate as task-driven entities capable of planning, interacting with compilers and debuggers, and self-correcting based on feedback.
Several trends motivated this change. First, real-world software development often requires iterative problem solving, tool use, and adaptation, which single-step code generation cannot handle effectively (liu2024large, ; li2024survey, ). Second, the rise of prompt engineering and structured prompting techniques (e.g., ReAct, chain-of-thought, scratchpads) enabled LLMs to reason more effectively over multiple steps (yao2023react, ; guo2025deepseek, ). Third, the increasing availability of APIs, command-line tools, and language server protocols made it possible to integrate LLMs into full-stack development environments (anthropic_claude_cli, ; google_cloud_cli, ; anthropic_mcp_spec, ).
These developments prompted a rethinking of how LLMs could be deployed-not just as smart coding assistants (github_copilot, ; tabnine, ), but as semi-autonomous agents capable of carrying out software engineering tasks with minimal human supervision. The resulting systems promise to augment developer productivity, lower the barrier to entry for software creation, and open new avenues for automation in programming workflows.
2.3. Agency in AI Systems
Agency is a foundational concept in the design of intelligent systems. At its core, an agent is an entity capable of perceiving its environment, reasoning about goals, and taking actions to influence outcomes. In the context of AI coding agents, agency refers to a system’s capacity to act autonomously, i.e., selecting actions based on internal objectives, external feedback, and learned knowledge.
Classical AI research has explored agency extensively in domains such as planning, robotics, and multi-agent systems (li2024survey, ; wu2024autogen, ). Traditional agent models focus on several key attributes:
  • Reactivity: the ability to respond to changes in the environment,
  • Proactivity: the pursuit of long-term goals,
  • Social ability: the capacity to communicate and coordinate with other agents or humans,
  • Autonomy: the ability to operate without direct human intervention.
In the context of AI agentic programming, agency manifests through an LLM-based system’s ability to:
  • Interpret open-ended tasks described in natural language,
  • Plan sequences of development steps (e.g., writing, testing, debugging),
  • Invoke and coordinate external tools (e.g., compilers, test runners, linters),
  • Adapt actions based on feedback from the environment (e.g., compiler errors or test failures),
  • Maintain coherent state and reasoning across multiple iterations.
Unlike symbolic AI agents that rely on explicitly defined world models and search-based planning, LLM-based coding agents operate in a probabilistic, language-driven manner. Despite this difference, they increasingly exhibit behaviors aligned with classical definitions of agency, especially when augmented with memory, tool-use modules, and planning routines.
2.3.1. Comparison with Robotics and Reinforcement Learning Agents
Robotic agents typically interact with the physical world through sensors and actuators. Their perception, control, and planning modules are tightly coupled with real-time feedback and often require safety guarantees. Reinforcement learning (RL) agents (sun2024llm, ), by contrast, learn behaviors by maximizing cumulative reward through trial and error, often in simulated environments. These agents explore large state-action spaces and learn policies over time.
Agentic programming systems share similarities with both paradigms:
  • Like robotic agents, coding agents must coordinate perception (e.g., task understanding) and action (e.g., code edits, tool invocations) in an environment with constraints and feedback.
  • Like RL agents, agentic systems benefit from feedback loops (e.g., test results, compiler outputs) and may incorporate exploration, retry strategies, or even reward-guided behavior.
However, coding agents operate in a symbolic, tool-rich environment, where actions are language-based, and environments (e.g., codebases, APIs, test harnesses) are highly structured. They must reason not only about immediate feedback but also about abstract software goals, dependencies, and long-term coherence across multiple steps. This makes agency in programming both uniquely challenging and distinctly different from physical-world or simulated agents.
2.4. Key Enablers of Agentic Programming
Figure 1 shows a representative system architecture of AI agentic programming. The emergence of AI agentic programming has been made possible by a combination of advances in language modeling, interaction frameworks, and software toolchains. Together, these enablers allow LLMs to move beyond static code generation toward goal-driven, interactive behavior. Below, we summarize the core technical factors that underpin this transition.
2.4.1. Large Language Models
Table 1. Representative LLMs for coding tasks.
Model Size (B) Context Win. Tool use Provider (access) Open Source MoE Used in coding IDEs GPT-5 N/A 1 M OpenAI (API only) VS Code, Cursor, other IDEs GPT-4 variants (o3, o4, etc.) N/A 128k OpenAI (API only) VS Code, JetBrains, Cursor Claude 4 Opus $\sim$300 200k Anthropic (API) Cursor, Replit (chat) Gemini 2.5 Pro $\sim$200 1M Google (API) Replit, Google Colab Grok 4 $\sim$1.7T $\sim$128k xAI Not publicly integrated DeepSeek R1-0528 671 (act. 37) 160k DeepSeek (API + weights) Emacs, VS Code (via extension) Kimi K2 1000 (act. 32) 128k Limited Moonshot AI (API) Custom plugin support Qwen3-235B-A22B 235 (act. 22) 32k Limited Alibaba Alibaba Cloud IDE Qwen3-Coder-480B-A35B-Instruct 480 (act. 35) 256k Alibaba Alibaba Cloud IDE Solar-Pro 72 128k Upstage (weights on HF) VS Code (via third-party) Openhands-LM-32B-v0.1 32 128k OpenHands VS Code (via extension) Devstral-Large 123 128k Mistral (API only) VS Code (via API) Devstral-Small 24 128k Mistral (API only) VS Code (via API)
LLMs trained on massive corpora of code and natural language form the foundation of modern agentic programming systems. These models, as represented by GPT-5 (openai2025gpt5introducing, ), Claude (claude2024, ), DeepSeek (guo2025deepseek, ), and Gemini (gemini2023, ), serve as the core reasoning engines, powering code generation, task planning, debugging, documentation, and natural language interaction. Their ability to understand and execute complex instructions makes them central to the design of agentic workflows.
Modern LLMs can generate syntactically correct and semantically meaningful code, answer development-related queries, and engage in multi-turn conversations with minimal task-specific fine-tuning. Many of these models leverage few-shot, zero-shot, and in-context learning capabilities, allowing them to generalize across programming languages, frameworks, and task domains. This flexibility enables developers to use the same underlying model for a wide range of software engineering tasks, from scaffolding and unit test generation to bug repair and performance tuning. In addition to general-purpose models, some LLMs like Grok (grok4, ) and Calude Opus (anthropic_opus4, ) are increasingly optimized for coding tasks through specialized instruction tuning, extended context length, tool use capabilities, and integration with retrieval-based systems. These enhancements make them suitable for multi-turn reasoning, code synthesis grounded in external context, and tool-augmented workflows.
Table 1 provides a comparative overview of state-of-the-art LLMs widely used in code-related tasks. The models listed include Grok 4, GPT-5 (GPT-4, o3 and o4 variants), Gemini 2.5 Pro, Claude 4 Opus, DeepSeek V3, Kimi K2, and Qwen3. The table compares their key attributes, such as the context window length. As the capabilities of LLMs continue to evolve, selecting and fine-tuning the appropriate foundation model becomes a critical design choice in building reliable, efficient, and adaptive agentic systems.
2.4.2. Prompt Engineering and Reasoning Strategies
Effective agentic behavior often requires structured prompting techniques to guide LLMs through multi-step reasoning and tool use. Methods such as chain-of-thought prompting, ReAct (reasoning and acting), scratchpad prompting, and modular prompting enable LLMs to plan, reflect, and revise outputs. These techniques allow agents to decompose complex problems, retain intermediate states, and act in ways that are more transparent and controllable.
2.4.3. Tool Use and API Integration
Table 2. Examples of tools supported by GitHub Copilot Agent.
Tool Type Examples Compiler gcc (GCC, ), clang (Clang, ), javac (Javac, ), tsc (TSC, ) Debugger gdb (GDB, ), lldb (LLDB, ), pdb (PDB, ) Test Framework pytest (Pytest, ), unittest (Unittest, ), Jest (Jest, ), Mocha (Mocha, ) Linter eslint (ESLint, ), flake8 (Flake8, ), black (Black, ), prettier (Prettier, ) Version Control git (Git, ) Build System make (Make, ), cmake (CMake, ), npm (NPM, ), maven (Maven, ) Package Manager pip (Pip, ) yarn (Yarn, ), cargo (Cargo, ) Language Server pyright (Pyright, ), tsserver (TSServer, )
Agentic systems rely heavily on external tools, such as compilers, debuggers, test frameworks, linters, and version control systems, to validate and refine generated code. Integration with these tools via command-line interfaces, language server protocols (LSP), or RESTful APIs enables agents to execute actions, gather feedback, and close the loop between generation and validation. Tool use is a defining feature of agentic programming, supporting iterative workflows and grounding decisions in observable outcomes.
For example, Table 2 lists some of the tools supported by GitHub Copilot Agent. The set of tools ranges from compilation to testing and version control.
2.4.4. State and Context Management
LLMs operate under fixed context windows, limiting their ability to reason over long histories. Agentic systems therefore incorporate external memory mechanisms to store plans, results, tool outputs, and partial progress. This memory can take the form of vector stores, scratchpads, or structured logs, allowing the agent to recall relevant information across multiple steps and maintain coherence over long-running tasks.
Table 3 compares the context management strategies of mainstream AI coding agents, revealing substantial differences in context size and memory persistence. Tools like GitHub Copilot and Codeium lack persistent memory, instead using transient methods such as sliding windows or dynamic token budgeting. In contrast, agents like SWE-agent, Devika, and OpenDevin employ persistent storage, often via vector databases or structured stores, to support long-term recall of plans, tool outputs, and project history. Some, such as Cursor IDE and Continue.dev, use embedding-based search to retrieve semantically relevant content, while others summarize prior actions to stay within the available context window.
These differences reflect a clear trade-off: smaller context windows typically rely on lightweight retrieval or summarization, whereas larger windows with persistent memory enable richer state tracking but add storage and retrieval overhead.
2.4.5. Feedback Loops and Self-Improvement
Agentic programming leverages feedback to refine outputs iteratively. Agents may rerun failed tests, revise prompts based on compiler errors, or reflect on past failures to improve future behavior. Some systems incorporate explicit planning, retry mechanisms, or even gradient-based updates through fine-tuning or reinforcement learning. This closed-loop design supports robustness and adaptability in complex programming tasks.
Together, these enablers mark a significant departure from traditional code generation pipelines. They allow LLMs to operate as interactive systems capable of planning, acting, and learning, paving the way for more autonomous and capable software development agents.
2.5. Comparison to Related Paradigms
AI agentic programming represents a distinct paradigm that builds upon but fundamentally differs from existing paradigms that have shaped the landscape of automated software development.
2.5.1. Program Synthesis
Program synthesis has been a foundational approach to automated code generation, traditionally divided into two types: deductive synthesis uses formal specifications to generate provably correct programs, while inductive synthesis learns from input-output examples with symbolic search and logic programming techniques to infer program logic (gulwani2017program, ; jain2022jigsaw, ). Classical synthesis systems like sketching (solar2008program, ) and more recent neural approaches like RobustFill (devlin2017robustfill, ) specialize in generating targeted code snippets that satisfy precise specifications.
However, program synthesis focuses on single-function generation from formal specifications and typically operates in a one-shot generation mode (devlin2017robustfill, ), whereas agentic programming handles multi-step workflows (e.g., planning, tool use, and iterative refinement) and engages in continuous interaction with development environments (bouzenia2024repairagent, ).
Table 3. Context management mechanisms supported by mainstream AI coding agents.
Agent Context Window Persistent Memory Context Management Mechanism GitHub Copilot 16k Sliding window over active buffer Codeium 32k Dynamic token budgeting based on file proximity and edit history Cursor IDE 128k Semantic search over project history SWE-agent 16k Vector DB retrieval for tool outputs and plan state Devika 32k Structured memory via SQLite and embeddings AutoDev 16k summarization of prior actions and tool logs Continue.dev 32k Embedding-based local recall over recent edits OpenDevin 32k RAG over command history, plans, and intermediate outputs
2.5.2. Code Completion Tools
Code completion, as one of the most commercially successful applications of AI in programming, excels at context-aware code suggestion, leveraging large-scale pre-training on code repositories to predict next tokens of partially written code (DBLP:conf/iclr/codegen23, ; DBLP:journals/corr/abs-2203-07814, ; husein2025large, ). Advanced completion tools can suggest entire functions, classes, or small modules based on comments, function signatures, and surrounding context, with tools like GitHub Copilot (github_copilot, ), TabNine (tabnine, ), and Amazon Q Developer (amazon-q-developer, ) achieving widespread adoption.
These code completion tools often operate as reactive assistants that respond to developer input, while agentic programming systems demonstrate proactive behavior and autonomous planning. Furthermore, agentic programming extends beyond code generation to encompass testing, debugging, deployment, and maintenance activities that completion tools typically do not address (yang2024swe, ; bouzenia2024repairagent, ).
2.5.3. DevOps Automation
DevOps automation focuses on streamlining software delivery pipelines through Infrastructure as Code (IaC), Continuous Integration/Continuous Deployment (CI/CD), and automated testing frameworks (humble2010continuous, ). Tools like Jenkins (jenkins, ), GitLab CI (gitlab_ci, ), and modern platforms like GitHub Actions (github_actions, ) automate repetitive deployment tasks, testing workflows, and infrastructure management.
While both paradigms emphasize automation, DevOps automation primarily handles pre-defined workflows and infrastructure management, whereas agentic programming focuses on adaptive problem-solving and creative solution generation. Additionally, agentic programming can potentially orchestrate and improve DevOps processes themselves, representing a higher-order form of automation (kambala2024intelligent, ).
2.5.4. AutoML and Automated Development
Automated Machine Learning (AutoML) represents a successful paradigm for democratizing AI model development through automation of model selection, hyperparameter tuning, and feature engineering (he2021automl, ). Platforms like Google Cloud AutoML (google_automl, ), Amazon SageMaker Autopilot (sagemaker_autopilot, ), and open-source frameworks like Auto-sklearn automate the traditional machine learning pipeline from data preprocessing to model deployment (feurer2015efficient, ).
However, AutoML focuses on statistical model optimization within well-defined machine learning workflows and operates with structured data and standardized evaluation metrics, while agentic programming tackles more general software development challenges with creative problem-solving and multi-modal reasoning.
2.5.5. Multi-Agent Systems and Human-AI Collaboration
Traditional multi-agent systems in software development typically involve specialized agent roles working within predefined coordination protocols (jennings2000agent, ). These systems often feature separate agents for requirements analysis, code generation, testing, and documentation, coordinating through structured communication interfaces. Recent advances in LLM-based multi-agent programming have demonstrated the potential for more sophisticated collaboration, with systems like MetaGPT showing how multiple AI agents can simulate software development teams (hong2023metagpt, ).
AI agentic programming can be viewed as an evolution of multi-agent systems that incorporates human-in-the-loop collaboration and dynamic role adaptation. Unlike traditional multi-agent systems with fixed agent roles and rigid communication protocols, agentic programming systems demonstrate fluid role assignment and context-adaptive behavior. Moreover, the integration of tool use, environmental interaction, and persistent memory distinguishes modern agentic programming from earlier multi-agent approaches. Contemporary agentic systems like AutoGen (wu2024autogen, ) and CrewAI (duan2024exploration, ) enable agents to directly interact with development tools, maintain context across extended sessions, and learn from past interactions (li2024survey, ). This represents a significant advancement over traditional multi-agent systems that typically operated in more constrained, simulation-based environments.
3. Survey Methodology
This survey follows a widely-used systematic literature review (SLR) methodology (keele2007guidelines, ; DBLP:journals/corr/Gong24Deep, ; hou2024largelanguagemodelssoftware, ; DBLP:journals/tse/WangHGGZFSLZN23, ; zhang2024systematic, ; qudus2025fact, ) to provide comprehensive coverage of AI agentic programming research, as illustrated in Figure 5.
3.1. Search Strategy
We conducted automatic searches across multiple academic databases, including Google Scholar, ACM Digital Library, IEEE Xplore, SpringerLink, and arXiv.org. We also examined proceedings from top-tier venues (FSE, ICSE, ASE, ICML, NeurIPS, AAAI, etc.).
Our search string combined the following term clusters using Boolean operators:
Agent terms: “AI agent” OR “agentic” OR “autonomous agent” OR “coding agent” OR “software agent” OR “intelligent agent” OR “task agent” OR “LLM agent” Programming terms: “programming” OR “coding” OR “software development” OR “code generation” OR “software engineering” OR “developer” OR “autonomous coding” OR “software automation” AI/LLM terms: “large language model” OR “LLM” OR “language model” OR “foundation model” OR “AI model” OR “neural code generation”
3.2. Study Selection
After initial retrieval, we followed a three-stage study selection process: (1) title and abstract screening by two independent researchers, (2) full-text review with disagreement resolution through discussion, and (3) backward and forward citation chaining to identify additional relevant studies. During the selection process, we used the following criteria:
Inclusion criteria - studies were included if they met all of the following:
  1. (1)
    Focus on AI systems for software development with autonomous/semi-autonomous behavior
  2. (2)
    Demonstrate agentic behaviors: planning, tool use, iterative refinement, or adaptive decision-making
  3. (3)
    Present novel techniques, architectures, evaluations, or comprehensive analysis
  4. (4)
    Include experimental evaluation, case studies, or substantial implementation details
  5. (5)
    Written in English with accessible full text
Exclusion criteria - studies were excluded if they:
  1. (1)
    Focused solely on traditional code completion without agentic behavior
  2. (2)
    Addressed non-programming domains (robotics, game playing, etc.)
3.3. Results
Our systematic search yielded:
  • Initial retrieval: 7,700 papers from database searches
  • Title/abstract screening: 395 papers selected for full-text review
  • Full-text review: 141 met all criteria.
  • Final corpus: 152 papers included after full-text evaluation and citation chaining
Among the 152 academic references published from 2022 to 2025 (excluding tool descriptions and websites), 5% appeared in 2022, 22% in 2023, 53% in 2024, and 20% in 2025, as shown in Figure 5, reflecting a surge in AI agent programming research following the widespread adoption of LLMs.
Figure 4. Survey Methodology.
[…]
[width=]figures/pie_year
Figure 5. Distribution of Papers.
4. Taxonomy of AI Agentic Programming
AI agentic programming is an emerging paradigm that equips LLM-based systems with autonomy, enabling them to plan, execute, and refine programming tasks over multiple steps. To provide structure to the diverse and fast-evolving landscape of agentic programming systems, this section introduces a taxonomy based on key behavioral and architectural dimensions. We categorize existing systems and approaches along these axes to clarify the design space and inform future development.
4.1. Agentic Behaviour Dimensions
This subsection defines the primary behavioural traits that differentiate agentic systems, forming the basis for a comparative classification.
4.1.1. Reactivity vs. Proactivity
Reactive agents respond directly to user prompts or feedback without independent task planning. For example, GitHub Copilot reacts by instantly suggesting a function body based on context after users type a function header like def initial. It does not include subsequent steps like writing tests or checking generated code. Proactive agents initiate sub-tasks, form execution plans, and re-evaluate decisions, often working autonomously over extended periods. For example, a proactive LLM-based agent can decompose a task into subtasks and execute them automatically. Users providing high-level instructions, such as “Add a user authentication module,” will generate the login logic, update the database, integrate it with the UI, and write corresponding unit tests.
4.1.2. Single-turn vs. Multi-turn Execution
Single-turn agents perform actions in response to individual prompts, often without preserving context. For example, classical GitHub Copilot responds to each prompt independently, without remembering past interactions. In contrast, multi-turn agents, such as GitHub Copilot Agent or Claude Opus 4 with tooling capabilities, maintain state across interactions, enabling iterative refinement, exploration, and goal pursuit. For instance, GitHub Copilot Agent can hold a conversation across multiple steps, remember earlier function names, and build a complete module through back-and-forth iterations between agents (github_copilot, ).
4.1.3. Tool-Augmented vs. Standalone Agents
Some agents are tightly integrated with external tools (e.g., compilers, debuggers, browsers, test frameworks), allowing them to perform code execution, validation, and correction. Others operate solely within the LLM’s reasoning capabilities, limiting their interactivity and adaptability.
4.1.4. Static vs. Adaptive Agents
Static agents (e.g., GitHub Copilot and Tabnine) follow predefined workflows or heuristics. Adaptive agents modify their strategies using feedback from tools, user input, or environmental signals. Some employ learning mechanisms to improve over time. For example, GitHub Copilot Agent adapts its approach when test failures occur, revising its implementation or replanning subtasks.
4.2. Agent System Categories
We now present a classification of current AI agentic programming systems, organised by their core functionality and architectural patterns.
4.2.1. Interactive Code Assistants
These are among the most widely adopted applications of LLMs in software development. These systems assist developers by providing code completions, inline documentation, editing suggestions, and simple refactorings. They are typically integrated directly into editors and IDEs, where developers interact with the underlying LLMs either through chat-like interfaces or by selecting code or comments using mouse-based interactions.
GitHub Copilot (github_copilot, ) and Cursor (cursor, ) are two representative examples of LLM-based code assistants. GitHub Copilot, originally developed based on Codex trained on GitHub code repositories, offers context-aware code completions across multiple programming languages and is tightly integrated into popular IDEs such as Visual Studio Code (vscode, ) and JetBrains. Cursor extends this functionality by embedding conversational interaction, maintaining memory of previous edits, and supporting structured command execution, allowing for a more dynamic and iterative user experience. Other notable systems include Amazon Q Develiper (amazon-q-developer, ) and Tabnine (tabnine, ). Amazon CodeWhisperer targets developers working in cloud ecosystems, offering language-specific completions, cloud API integration, and basic vulnerability detection. In contrast to CodeWhisperer, Tabnine takes a privacy-first approach by deploying smaller local models trained on permissively licensed code, making it attractive for organizations and developers who do not want to send their code to a remote cloud.
Implementations of these systems typically exhibit reactive behavior, responding to user input without initiating their own plans or taking proactive steps. Their interactions are generally single-turn, relying on the immediate context within the code editor rather than maintaining a persistent memory of past interactions or broader development goals. Most systems in this category are tightly coupled with development tools and offer real-time assistance that fits naturally within existing programming workflows. However, they are limited in autonomy, lacking the ability to decompose complex tasks, maintain long-term state, or coordinate multi-step development activities.
Despite these limitations, interactive code assistants serve as a foundational layer within the more recent agentic programming ecosystem. They are widely deployed, easy to integrate into everyday development practices, and offer immediate value to developers. Furthermore, some recent systems, such as Cursor and GitHub Copilot Agent, are beginning to incorporate features like session-level memory, persistent context, and structured task execution, gradually bridging the gap between reactive code assistants and more autonomous, multi-turn agents.
These distinctions are captured in our taxonomy (see Table 4), where interactive code assistants are compared with other categories of agentic systems across key dimensions, including autonomy, memory scope, tool integration, reasoning complexity, and interaction model.
4.2.2. Autonomous Task-Oriented Agents
These agents perform multi-step programming tasks with minimal human intervention, often maintaining control over the entire development process from requirement interpretation to code generation and validation. They can plan, execute, and revise their own workflows in response to intermediate results or changing task requirements. Many integrate external tools such as debuggers, package managers, or search engines, enabling them to gather information, resolve errors, and optimize code without direct user guidance. Examples include the state-of-the-art LLMs that have tool integration capabilities like GPT-5 (openai2025gpt5introducing, ), Claude Opus 4 (anthropic_opus4, ), Google Jules (Jules, ), Kimi K2 (team2025kimi, ), and Gemini 2 (comanici2025gemini25pushingfrontier, ). These agents are often proactive in suggesting next steps, adaptive to new inputs, and capable of maintaining continuity across extended sessions through persistent memory or context management mechanisms.
4.2.3. Planning-Centric Agents
Planning-centric agents approach problem-solving as a two-phase process: first, structured task decomposition, where high-level goals are broken down into smaller, more manageable steps; and second, execution monitoring, in which results are evaluated and the plan is adjusted accordingly. This approach improves the handling of long-horizon tasks. For example, CAMEL (li2023camel, ) employs two agents in a role-playing setup, a “user” agent and an “assistant” agent, which collaboratively refine goals and strategies, producing plans that downstream code generators can execute. These agents are typically multi-turn, memory-enabled, and trading speed for robustness.
4.2.4. Multi-Agent and Collaborative Systems
Multi-agent and collaborative systems extend agent-based programming by introducing multiple specialized agents that coordinate to solve complex software engineering tasks. This approach draws inspiration from human software teams, where each member has a distinct role, such as requirements analysis, coding, testing, or documentation, and communication protocols are established to ensure progress toward shared objectives. For example, SWE-Agent (yang2024swe, ) employs multiple role-specific LLM agents—an “Architect” agent for high-level design, a “Coder” agent for implementation, and a “Reviewer” agent for quality assurance—connected through structured dialogue and shared memory. Such systems offer high autonomy, scalability, and tool integration, but face challenges in avoiding redundancy and managing conflicts.
4.3. Summary and Comparative Table
We conclude this section with a comparative summary (Table 4) of representative systems across the behavioural dimensions and categories introduced above. This taxonomy provides a lens to understand the capabilities and limitations of existing approaches and can guide the design of future systems.
Table 4. Comparison of Representative AI Agentic Programming Systems.
System Category Proactivity Multi-turn Tool Use Adaptivity
GitHub Copilot IDE Assistant Reactive
Tabnine IDE Assistant Reactive
Cursor IDE Assistant Reactive
Claude Opus 4 Task-oriented Proactive
Google Jules Task-oriented Proactive
Dev-GPT Task-oriented Proactive
KimI K2 Task-oriented Proactive
Gemini 2 Task-oriented Proactive
Voyager Planning Agent Proactive
CAMEL Planning Agent Proactive
CodePlan Planning Agent Proactive
ChatDev Multi-agent System Proactive
AutoCodeRover Multi-agent System Proactive
SWE-Agent Multi-agent System Proactive
4.4. Cost and Token Consumption Model
While recent LLMs show impressive capabilities in software engineering tasks, their real-world applicability is often constrained by cost considerations. This cost is typically measured in terms of tokens consumed per US dollar for both input and output, along with additional expenses incurred by extended reasoning strategies such as Chain-of-Thought (CoT) and tool-augmented workflows.
4.4.1. Pricing Dimensions.
Commercial providers generally price LLM usage by input and output tokens, with rates varying across model families. Some, such as OpenAI’s GPT-5, offer multiple service tiers (Standard, Mini, Nano, Pro) with different pricing, context lengths, and throughput limits. Table 5 summarizes representative pricing.
Table 5. Representative token pricing for LLMs used in coding tasks (USD per 1M tokens, as of Aug. 2025).
Model Input(\$) Output(\$) Context Window
GPT-5 (Standard) 1.25 10.00 256k
GPT-5 Mini 0.25 2.00 128k
GPT-4 variants (e.g., 4o) 2.50 10.00 128k
Claude 4 Opus 15.00 75.00 200k
Gemini 2.5 Pro 1.25 10 1M
Grok 4 3.00 15.00 256k
DeepSeek R1-0528 0.55 2.19 160k
Kimi K2 0.15 2.50 128k
Qwen3-235B-A22B 0.22 0.88 32k
Qwen3-Coder-480B-A35B-Instruct 4.5 13.5 256k
Solar-Pro 0.30 0.30 128k
Openhands-LM-32B-v0.1 2.6 3.4 128k
Devstral-Small (Mistral) 0.10 0.30 128k
Devstral-Large (Mistral Medium 3) 0.40 2.00 128k
4.4.2. Impact of Reasoning Strategies.
To make the cost-performance trade-offs more concrete, we draw on the “Agentic workflow for implementing a REST task” example described earlier in Section 2. In that workflow, the agent receives a high-level natural language specification for a REST API endpoint, interprets the requirements, generates code, and iteratively tests the implementation until it passes the provided unit tests.
We consider three reasoning strategies applied to this scenario:
Short reasoning. The model produces the implementation in a single turn with minimal intermediate reasoning. For the REST API task, this means directly generating the endpoint code and tests without explicit planning or validation steps. This approach minimizes token usage and latency but risks missing subtle requirements.
Standard CoT. The model uses a fixed-depth chain of thought to plan the implementation. In the REST API case, this involves reasoning about request handling, data validation, and error responses before generating the code. This strategy consumes significantly more tokens, roughly twice as many, compared to the short reasoning strategy, but yields a higher likelihood of producing a correct implementation on the first attempt.
Tool-augmented iterative reasoning. The agent integrates code compilation and test execution into the workflow. After producing an initial version, it runs the tests, inspects any failures, and revises the code in subsequent turns. For the REST API example, this may involve multiple cycles of fixing logic errors, adjusting request parsing, and refining edge-case handling. While this maximizes accuracy and robustness, it also increases token consumption and wall-clock time substantially due to repeated code generation and analysis.
In practice, the optimal choice depends on the cost-performance budget of the project. For time-sensitive or budget-constrained environments, a hybrid approach can offer a more effective balance.
5. Challenges
AI agentic programming introduces a promising and complex shift in how software is developed, relying on the autonomous capabilities of LLMs. Despite recent progress, several technical and conceptual challenges remain that hinder the deployment of robust, scalable, and trustworthy agentic systems (sapkota2025aiagentsvsagentic, ; acharya2025agentic, ).
5.1. Inadequate Evaluation and Benchmarking
A variety of benchmarks and open-source toolkits (liu2023agentbench, ; qin2023toolllmfacilitatinglargelanguage, ) have been proposed to evaluate the capabilities of LLM-based agents across different tasks. Despite this progress, existing benchmarks commonly used for assessing coding agents, such as HumanEval (zheng2023codegeex, ) and SWE-Bench (jimenez2023swe, ), remain inadequate for capturing the full complexity of real-world software engineering workflows.
These benchmarks primarily focus on small, self-contained problems, often restricted to a narrow range of programming languages, such as Python (chen2021evaluatinglargelanguagemodels, ), and generally lack support for interactive, multi-turn, or tool-integrated tasks (liu2023agentbench, ). In contrast, practical agentic systems are expected to operate over large, modular codebases, interface with third-party libraries, manage build workflow pipelines, and respond dynamically to user feedback or runtime tool outputs. As LLM-based systems increasingly incorporate reinforcement learning (ouyang2022training, ; wang2024survey, ) and more advanced planning mechanisms, future benchmarks should reflect this integration.
Moreover, there is a noticeable absence of evaluation frameworks designed for emerging complex use cases, such as those involving interactions with compilers and debuggers (wang2024executablecodeactionselicit, ), where agents must reason about low-level program behavior, perform iterative transformations, or track state across toolchains. The lack of such domain-specific benchmarks presents a significant gap in evaluating agent performance under realistic conditions.
5.2. Domain Foundation Models for Agents
Generic coding agents often struggle in domain-specific environments such as embedded systems, high-performance computing, optimization, or formal software verification (lecong2025llmsreasonprogramsemantics, ; loughridge2024dafnybench, ). These domains typically impose stricter operational constraints and require deep integration with specialized APIs, toolchains, and domain knowledge resources that are often underrepresented in general-purpose training corpora.
To address these limitations, recent research has proposed domain-adapted models and task-specific learning strategies to accelerate agent performance in specialized settings (song2025injecting, ). For instance, some approaches have begun incorporating compiler knowledge or security-specific patterns into LLM training pipelines (lin2025llmssecuritycoursesecuring, ; LLMCompiler, ), enabling agents to reason more effectively about low-level program behavior or vulnerability patterns.
In the future, developing robust and adaptable domain foundation models will be a promising direction for enabling agents to operate reliably in complex software environments, such as LLMs pretrained or fine-tuned on domain-specific data, tools, and semantics.
5.3. Safety and Privacy
As agentic systems gain increasing autonomy, so does the potential for unsafe behavior. Unlike traditional tools (Checkmarx, ), agentic systems can invoke external tools, perform structural code modifications, and even commit changes without direct human oversight (triedman2025multiagentsystemsexecutearbitrary, ; zhang2025agentsafetybenchevaluatingsafetyllm, ). These capabilities introduce significant risks, including the possibility of introducing subtle bugs, propagating unsafe patterns, or violating security constraints.
A critical future direction involves ensuring that agentic systems can protect users and data. For example, when agents visit private repositories or are deployed in cloud-integrated environments, future models may need built-in controls to restrict access to sensitive project data.
Further, malicious prompts, poisoned APIs, or compromised toolchains can mislead agents into executing unsafe behaviors (debenedetti2024agentdojodynamicenvironmentevaluate, ; zhan2024injecagentbenchmarkingindirectprompt, ). Future research should prioritize the design of secure protocols for agent collaboration, including authentication between agents, validation of tool outputs, and detection of anomalous actions.
Also, agents must be capable of explaining their reasoning, flagging uncertainties, and allowing developers to understand and revise with minimal effort. Building safety and privacy into the foundation of agentic architectures is essential.
5.4. Toolchain Integration and Programming Language Design
One fundamental challenge lies in the incompatibility between existing software tools and the needs of autonomous agents. Most programming languages, compilers, debuggers, and development tools are designed with human developers in mind (rorseth2025ladybug, ).
They emphasize usability over feedback, so agents often struggle to diagnose failures, trace the effects of their code transformations, or understand build errors (kim2024llm, ). For example, compilers typically provide minimal insights into transformation failures or semantic conflicts (li2025review, ).
To support agentic workflows, toolchains must evolve to expose IRs, transformation traces, and structured feedback interfaces. Moreover, language-level annotations and agent-aware interfaces are needed to communicate developer intent and guide automated reasoning.
5.5. Scalable Memory
Agentic programming systems must maintain coherence and reasoning over long-running tasks involving multiple iterations, tools, and contextual dependencies. However, current LLMs are limited by fixed context windows and lack persistent, structured memory mechanisms (wang2023augmenting, ).
Realistic software tasks, such as feature implementation, debugging, or refactoring, require agents to store and reason over evolving states, feedback logs, intermediate plans, and prior actions (qian2025memorag, ). Without hierarchical and queryable memory systems, agents risk repeating errors, forgetting past successes, or producing inconsistent results.
Emerging solutions such as retrieval-augmented generation and memory summarization offer partial relief, but they remain inadequate for complex, multi-session workflows (maharana2024evaluating, ). Future research can explore memory architectures that differentiate short-term interactions, mid-term subgoals, and long-term domain knowledge.
6. Opportunities and Future Directions
AI agentic programming represents a fast-evolving research frontier that intersects artificial intelligence, programming languages, and software engineering. While recent advances have demonstrated promising capabilities, significant challenges remain in realizing robust, efficient, and trustworthy agentic systems. In this section, we outline several key opportunities and open research directions that can shape the future of this field.
6.1. Integrating Coding Agents with Tools
Existing AI coding agents typically orchestrate LLMs with loosely integrated toolchains and basic memory mechanisms. These ad hoc designs often lack robustness, scalability, and generalization across programming tasks. Advancing agent architectures will require moving beyond simple prompt-response patterns toward more modular, structured systems that support reasoning, tool interaction, planning, and verification.
A promising direction is to rethink how programming languages, compilers, and testing frameworks, which are traditionally built for human developers, can be redesigned to support AI coding agents. For example, instead of emitting opaque diagnostics, compilers could provide structured feedback explaining why certain optimizations (e.g., vectorization or inlining) fail (maleki2011evaluation, ; ashouri2018survey, ; siso2019evaluating, ). These could include semantic barriers like unresolved aliasing, ambiguous data/control flow, or missing annotations (bi2024iterative, ), enabling agents to revise code more precisely. Beyond diagnostics, compilers can help agents track state across iterations. Feedback on which edits introduced errors, failed assertions, or performance regressions would enable agents to reason over the change history and adjust strategies accordingly.
Opening compiler internals also presents a valuable opportunity. Coding agents could interact directly with intermediate representations (IRs), such as LLVM IR (lattner2004llvm, ) or MLIR (lattner2021mlir, ), to reason about program structure, verify transformations, or perform static analysis at a semantic level. Compiler APIs and language servers (e.g., Clang’s LibTooling, the Language Server Protocol) already expose ASTs, symbol tables, and refactoring tools, but a wider adoption may require standardized, introspective interfaces across compilers.
At the programming language level, agent-aware extensions or annotations could further improve interaction. Developers might use domain-specific languages, embedded contracts, or even natural language comments, e.g., “sort the elements of input array xx”, to convey intent. This could guide synthesis, verification, or debugging. Likewise, compilers might expose symbolic summaries of control flow, memory access patterns, or performance profiles to inform multi-step agent reasoning.
Tighter integration with runtime systems also offers opportunities. For instance, agents can dynamically insert instrumentation or launch profiling runs, then use the results to inform optimization choices. Coupling these capabilities with autotuning frameworks (chen2018tvm, ; ansel2014opentuner, ) would expand the design space while preserving correctness and safety.
Finally, advances in structured code representations, such as ASTs, graph-based IRs, and semantic embeddings, offer a foundation for more powerful agent reasoning. Combining LLMs with graph neural networks or neuro-symbolic systems could improve generalization and support cross-language, cross-target understanding.
6.2. Scalable Memory and Context Management
A key capability of agentic programming lies in managing memory and contextual information across tasks that involve long context reasoning and multiple iterations. Unlike traditional code generation, which typically follows a single pass prompt to solution model, agentic workflows for solving real-world software engineering problems involve multiple steps, iterative refinement, and integration with external tools and development environments (cursor, ; wang2024openhands, ; openai2021codex, ; github_copilot, ).
Consider an agent tasked with adding a new feature to an open source project, such as implementing a command-line flag to enable verbose logging. The agent must first analyze the existing codebase to locate the argument parsing logic, generate the required code changes, and update the logging behavior. If the updated code fails with a runtime error due to an uninitialized flag, the agent needs to debug the issue by inspecting stack traces, revise the code accordingly, and rerun the tests. Once the implementation is verified, the agent writes a commit message, creates a pull request with a summary of the changes, and links it to the relevant issue.
Throughout this process, the agent must persist and reason over a large and evolving context: the initial task description, previously generated code, compiler and runtime feedback, and version control metadata. Without the ability to store and recall this information in a structured way, the agent may repeat past mistakes, forget earlier successful changes, or submit incomplete solutions. Agentic programming, therefore, needs mechanisms for memory and context tracking that go beyond simple token limits, enabling coding agents to maintain continuity across extended interactions and tool usage.
As the memory footprint of LLMs grows linearly with input token length (DBLP:conf/nips/VaswaniSPUJGKP17, ; kaplan2020scaling, ), current LLM-based agents remain constrained by their context windows and lack persistent memory across a long sequence of iterations (schick2023toolformer, ; liu2023lost, ; packer2023memgpt, ). However, reasoning about real-world programs often requires modeling complex data structures and code context (like function calls) spanning across multiple files, which often exceeds typical context limits. Although some industry-scale LLMs claim to support million-token contexts (ding2023longnet, ; team2024gemini, ; ding2024longrope, ), they often rely on random sampling techniques (hosseini2024efficient, ) and fail to leverage program structure or semantics effectively.
Therefore, an interesting direction is to design attention mechanisms that are guided by code structure, such as syntax trees, control flow graphs, or data dependencies. These structures can help agents focus more accurately on the most relevant parts of a program. While approaches like retrieval-augmented generation (RAG) (lewis2020retrieval, ), KV cache offloading (lee2024infinigen, ; sun2024shadowkv, ), and compression (liu2024kivi, ; chang2025palu, ) provide partial solutions, they struggle to provide precise control over long-term dependencies, structured knowledge, and execution histories.
AI coding agents can also benefit from hierarchical memory models that distinguish between short-term interaction history, mid-term planning objectives (such as subgoals and intermediate decisions), and long-term knowledge. This long-term layer may include patterns of success or failure, reusable code templates, and observed tool behaviors. Such hierarchies can be dynamically updated and selectively queried using retrieval controllers or attention-based mechanisms. Additionally, memory summarization techniques could be explored to condense lengthy interaction histories into structured, semantically meaningful representations. For example, an agent might summarize a multi-turn session as a sequence of planning decisions and outcomes, highlighting key insights and interventions.
Another important area is the development of context-aware retrieval strategies that move beyond static similarity-based methods (su2024dragin, ). During debugging, for instance, an agent could retrieve not only the most recent error message but also similar past failures, proposed fixes, relevant test cases, and their outcomes. Retrieval conditioned on task state and tool feedback would significantly improve the agent’s ability to reason under uncertainty.
Structured mechanisms for program state tracing and replay may also enhance agent performance. By recording partial program states, tool outputs, and execution steps, agents can support backtracking, recovery from failure, and richer explanations. For example, an agent could explain how a specific code edit introduced a type error or why a particular memory access blocked loop vectorization. These capabilities are crucial for supporting causal reasoning and improving transparency. Likewise, persistent memory across multiple code generation and refinement sessions will be essential for enabling agents to accumulate and refine knowledge over time. This may include long-term storage of project-specific context, interaction histories, usage patterns of tools, and models of user intent. Such memory infrastructure will allow for continual learning and increasing personalization of agent behavior.
In summary, effective memory and context management are foundational for scaling agentic programming systems. These capabilities are vital for advancing from reactive, prompt-driven assistants to autonomous, context-aware collaborators capable of sustained reasoning, adaptation, and long-term learning.
6.3. Evaluation and Benchmarking
Table 6. Benchmarks for Evaluating LLMs and Agentic Systems on Programming Tasks.
Abbreviations: CP = Competitive Programming, BF = Bug Fixing, FC = Function Completion, CR = CLI Reasoning.
Benchmark Source Language Task Difficulty Year
HumanEval Hand-written Python FC Beginner 2021
MBPP Crowd-sourced Python FC Beginner 2021
CodeContests CP Python/C++/Java FC Diverse 2022
HumanEval-X Hand-written Python/C++/Java/JS/Go FC Intermediate 2023
SWE-Bench GitHub Issues Python BF Expert 2024
SWE-bench M GitHub Issues JS BF Diverse 2024
LiveCodeBench CP (live) Python FC, BF Diverse 2024
TerminalBench Community-curated Shell CR Diverse 2025
Spider 2.0 Enterprise DB Apps SQL FC Expert 2025
EffiBench-X Synthetic Python/C++/JS FC, BF Diverse 2025
Web-Bench Web App Projects JS/TS/HTML/CSS/Python BF, FC Expert 2025
ProjectEval Open-source repos Python/Java/C++/JS FC, BF, CR Expert 2025
TRAIL CP Python/Java/C++/JS BF, CR Expert 2025
Note: ”Diverse” difficulty indicates that the benchmark covers a wide range from beginner to expert-level tasks.
Table 7. AI Intelligence Score (across mathematics, science, coding, and reasoning) and Speed (Output tokens per second) for LLMs on LivecodeBench
Metric GPT-5 GPT-4o Claude 4 Gemini 2.5 Pro Grok 4 DeepSeek R1 Kimi K2 Qwen3 Solar-Pro Devstral Llama 4 Magistral
AI score 69 65 59 65 68 59 49 64 43 31 42 36
Speed 102 148 100 153 76 21 45.2 24 54 44 174 51
Table 6 summarizes several widely used coding benchmarks for evaluating LLMs and agentic systems on programming tasks. These benchmarks vary in their origin, programming language coverage, task types, and difficulty levels. Commonly used datasets include HumanEval (openai2021codex, ), HumanEval-X (zheng2023codegeex, ), MBPP (austin2021program, ), SWE-Bench (jimenez2023swe, ), SWE-bench Multimodal (yang2024swe, ), TerminalBench (tbench_2025, ), LiveCodeBench (jain2024livecodebench, ), CodeContests (DBLP:journals/corr/abs-2203-07814, ), Spider2.0 (lei2024spider, ), EffiBench-X (qing2025, ), Web-Bench (xu2025webbench, ), ProjectEval (liu-etal-2025-projecteval, ), and TRAIL (deshpande2025trailtracereasoningagentic, ).
While these benchmarks have provided valuable insights into the capabilities of LLMs and agentic systems for code generation and bug fixing, they also have important limitations. For example, they are heavily biased toward a small set of programming languages - especially Python, which dominates the training corpus of code for current models (twist2025llms, ; chai2024mceval, ). This limits the generalizability of evaluation results to domains involving statically typed or domain-specific languages, such as C++ or Rust (twist2025llms, ). For example, Table 7 presents the AI Intelligence Score (aggregated across mathematics, science, coding, and reasoning) and the output speed (in tokens per second) for a selection of widely used LLMs on the LiveCodeBench evaluation. For the intelligence score, the latest model, GPT-5, achieved 69, indicating that more than 30% of challenges remain unresolved.
Furthermore, many existing benchmarks focus on small, self-contained problems that may not be representative of real-world software engineering tasks. Realistic development scenarios often involve working with large, modular codebases, extensive use of third-party libraries, non-trivial build processes, and long-range dependencies across files and components. These characteristics are largely absent from current benchmark datasets, making it difficult to assess an agent’s ability to scale or generalize.
Another key limitation is that most benchmarks do not capture the interactive, iterative nature of agentic programming. In real-world settings, coding agents will need to collaborate with human developers (wang2024openhands, ; cursor, ; openai2021codex, ; github_copilot, ), receive intermediate feedback and confirmation, and rely on external tools such as compilers, debuggers, and test frameworks. Benchmarks that assume single-shot or non-interactive task completion fail to reflect the complexity of such multi-step, tool-augmented workflows.
Addressing these gaps will require the development of more comprehensive and extensible evaluation frameworks. Future benchmarks should incorporate realistic tasks that reflect end-to-end development workflows, support multiple programming languages, and enable interaction with tools and human feedback loops. Metrics should go beyond functional correctness to include robustness, tool usage efficiency, recovery from failure, and the ability to incorporate feedback. Simulation environments and evaluation harnesses will also be important for reproducibility and fair comparison.
6.4. Human-AI Collaboration
While a long-term vision of agentic programming is to automate the entire software development lifecycle, including writing, debugging, and testing code, near-term opportunities include extending the capabilities of current LLM-based coding assistants (wang2024openhands, ; cursor, ; openai2021codex, ; github_copilot, ). Rather than replacing human developers, these systems can act as collaborative partners, supporting workflows in which humans retain strategic oversight. Similar to pair programming (hannay2009effectiveness, ), LLM-based agents can assist by proposing ideas, detecting errors, suggesting improvements, and automating routine tasks, thereby augmenting human productivity.
Designing effective human–AI collaboration models remains an important research challenge. This includes developing user interfaces, interaction protocols, and frameworks for allocating responsibilities between human developers and agents, particularly in professional and team-based environments where coordination, trust, and efficiency are essential. Unlike traditional code generation tools, agentic systems operate in iterative, feedback-driven loops and are capable of making autonomous decisions based on intermediate results. This interaction model shifts the role of the developer from directive command-giver to interactive collaborator, opening up new opportunities and challenges for co-creating software. To enable meaningful collaboration, agents must be transparent in their reasoning, responsive to human input, and able to explain their decisions. Realizing these capabilities will require advances in interactive prompting (schulhoff2024prompt, ), natural language explanations (creswell2022faithful, ), and context-aware dialogue protocols (shah2025prompt, ; anthropic2024mcp, ).
One promising research direction is the development of mixed-initiative workflows, in which control shifts fluidly between the developer and the agent. In these workflows, the human might specify high-level goals, architectural constraints, or coding conventions, while the agent proactively generates scaffolding code, explores design alternatives, or automates repetitive subtasks. The developer can then inspect, accept, reject, or refine the agent’s contributions. For instance, an agent could propose multiple refactoring strategies for a large function and generate test cases to verify the correctness of each variant. The human collaborator evaluates these options and selects the most appropriate one, potentially with further refinements. Supporting such workflows requires robust management of shared context, memory, and goal state across multiple interaction turns.
Agents must also be able to handle ambiguity and uncertainty. Real-world development often involves vague or underspecified requirements. Future research may explore agents that can ask clarifying questions (zhang2024modeling, ; mu2024clarifygpt, ), infer intent from surrounding artifacts, or learn user preferences over time. For example, when given an instruction like “make this more efficient”, the agent should determine whether to optimize for speed, memory, or readability - potentially by prompting the user for clarification.
Furthermore, human developers rely on a rich ecosystem of non-code artifacts, such as design diagrams, issue trackers, documentation, and domain-specific knowledge. Integrating these contextual signals into the agent’s reasoning pipeline remains an open challenge. For instance, when fixing a bug, the agent may need to cross-reference issue tracker discussions or past commits to fully understand the context and constraints of the fix.
Trust and usability are foundational to collaboration. Developers must feel confident that the agent’s actions are safe, interpretable, and reversible. Research is needed on techniques that allow agents to explain their decisions, cite relevant documentation or examples (e.g., from internal codebases or public repositories), and highlight trade-offs. Users should be empowered to inspect, override, or revert agent-generated changes easily—especially in safety-critical or compliance-sensitive domains such as finance, healthcare, or aerospace.
Collaboration in team-based environments presents additional opportunities. Future systems may support multi-user workflows where agents assist in coordinating tasks across a team, help onboard new developers by answering project-specific questions, or even mediate merge conflicts by suggesting consistent resolutions based on prior history. For example, an agent integrated into a version control system might suggest commit messages, flag inconsistent code style across pull requests, or recommend reviewers based on code ownership.
In summary, enabling rich human-AI collaboration in agentic programming will require progress across multiple dimensions: user interface design, communication strategies, intent modeling, memory management, and trust calibration. These systems should be designed not to replace human developers, but to augment their creativity, efficiency, and decision-making. Ultimately, effective collaboration will depend on how well agents can adapt to the social and cognitive workflows of modern software development.
6.5. Domain Specialization and Adaptability
General-purpose coding agents often underperform in domain-specific tasks due to a lack of tailored knowledge, tool integration, and contextual understanding (joel2024survey, ; zheng2024well, ). While LLMs trained on extensive code corpora can demonstrate general programming capabilities, they may struggle in specialized domains, such as generating Verilog code for hardware synthesis, which involves strict performance constraints, low-level abstractions, or domain-specific libraries and tooling.
Future research may explore strategies for domain-adaptive prompting, fine-tuning, and agent behavior modulation to improve performance in areas such as embedded systems, data science, scientific computing, high-performance computing, and formal methods. For example, in embedded systems programming (englhardt2024exploring, ), agents must reason about memory layout, real-time constraints, and hardware-specific APIs, which are often underrepresented in general training datasets of code. Fine-tuning on domain-specific codebases or adapting the agent’s planning strategy to prioritize correctness and safety over code brevity may significantly improve performance. Another direction is to build LLM agents that incorporate domain-specific tools and diagnostics into their reasoning loop. For instance, a scientific computing agent might integrate with tools like NumPy, MATLAB, or domain-specific simulation engines (openfoam, ; permann2020moose, ; simulink, ), allowing it to generate, run, and analyze numerical experiments autonomously. Similarly, an agent for cybersecurity tasks might interact with static analyzers (codeql2020, ), fuzzers (manes2019art, ), or formal verification tools (woodcock2009formal, ) to check code properties as part of its development process.
Domain specialization can also improve interpretability and safety. By narrowing the operational scope of an agent, developers can apply stricter validation, leverage existing formal specifications, and generate more meaningful feedback. In safety-critical applications, such as aerospace, automotive, or medical software, agents must operate under rigorous constraints, and tailoring their reasoning and code generation behaviors to domain-specific requirements is essential.
Additionally, research into adaptive agent behavior can enable models to detect the domain of a task at runtime and adjust their prompting strategies, tool usage, and explanation styles accordingly. For example, an agent working on a financial modeling script might switch to using Pandas and SQL queries, while one dealing with real-time control code may emphasize low-latency function design and interrupt safety. This line of research will also open opportunities for personalisation at the developer or team level. Agents can learn preferences, coding conventions, and domain-specific heuristics from local repositories or past interactions, enabling more consistent and context-aware assistance.
6.6. Safety, Alignment, and Trust
As agentic coding systems become more autonomous and start taking on greater responsibility in software development, it becomes increasingly important to ensure that their behavior aligns with user intent, produces correct results, and avoids unintended changes (ouyang2022training, ; liu2024exploring, ). Unlike traditional LLM-based code assistants, agentic systems can take multi-step actions, use external tools, and modify codebases with limited human oversight, making safety and trust essential goals (nvidia2025safetyrecipe, ; hua2024trustagent, ).
One key research direction is building agents that can better understand and follow what users actually want, even when instructions are vague or incomplete. Current systems rely heavily on natural language prompts, which can be ambiguous. Future work could focus on grounding agent behavior in more structured forms of input, such as constraints, test cases, or high-level goals, that are easier to validate and reason about (fakhoury2024llm, ; huang2023agentcoder, ; li2025structured, ).
An interesting idea is to design a structured language that developers can use to express their intent clearly. This language would serve as a kind of programming interface for LLMs, allowing users to define what the agent should do, what it must avoid, and what counts as a valid solution. For example, a developer might specify that a function’s output must remain the same after refactoring, and the agent would only explore changes that meet this requirement. This kind of structure could also make it easier to apply lightweight verification tools, such as static analyzers or type checkers, to ensure the agent’s suggestions are safe and correct (codeql2020, ; li2024iris, ; nunez2024autosafecoder, ).
In safety-critical domains like healthcare, finance, or automotive software, agents will also need to follow strict coding standards, legal rules, and certification guidelines. Agents could be trained to recognize such constraints or be paired with rule-based systems to flag violations and suggest compliant alternatives (ouyang2022training, ; he2024instruction, ).
Another important challenge is ethical alignment. Because agentic systems are trained on large and diverse codebases, they may learn unsafe or biased practices (liu2023uncovering, ; pearce2025asleep, ; huang2024bias, ). Research efforts are needed to detect and reduce risks such as generating insecure code, leaking sensitive data, or reinforcing stereotypes. Techniques from responsible AI, like behavioral audits, adversarial testing, and human feedback, can be adapted to this setting.
To build user trust, agents should also be able to explain their decisions (rajani2019explain, ; creswell2022faithful, ; DBLP:conf/nips/ShinnCGNY23, ). Developers need to know why an agent made a change, where the idea came from, and what trade-offs are involved. Research on explanation generation, source citation, and visualization of code changes will help make agent behavior more transparent and understandable (liu2024reliability, ; sun2022investigating, ; huang2023citation, ).
Agents should also be aware of their own limitations. When they are uncertain, they should be able to flag their confidence level (kuhn2023semantic, ; creswell2022faithful, ), suggest multiple options (DBLP:conf/nips/MadaanTGHGW0DPY23, ; DBLP:conf/nips/ShinnCGNY23, ), or ask the user for confirmation (mu2024clarifygpt, ; zhang2024modeling, ). Designing agents that can adapt their behavior based on task difficulty or user feedback is a promising direction (DBLP:conf/nips/ShinnCGNY23, ; sun2023adaplanner, ; damani2024learning, ; manvi2024adaptive, ; chen2024magicore, ).
Finally, safety mechanisms like undo and audit trails will be essential. Agents should keep track of what they have changed and allow users to roll back actions easily (patil2024goex, ; jiang2024rocode, ). Future work could explore automatic snapshotting, reversible code edits, and tight integration with version control systems to support safe collaboration.
7. Conclusion
We have presented a comprehensive review of AI agentic programming. This new software paradigm, driven by the success of large language models (LLMs), is a transformative shift in how software can be created, maintained, and evolved. By combining the capabilities of LLMs with planning, tool use, and iterative refinement, coding agents are beginning to automate complex, multi-stage programming workflows that traditionally required significant human involvement. These systems can not only generate and test code but also interact with development tools, decompose tasks, and adapt based on feedback, bringing us closer to the vision of autonomous software development.
In this article, we have introduced a taxonomy of AI coding agents and architectures, reviewed underpinning techniques like context management and tool integration, and discussed how existing benchmarks evaluate the capabilities of coding agents. We have summarized the progress made in enabling LLM-based AI agents to reason over tasks, interface with software development tools, and operate in increasingly sophisticated ways. At the same time, we have identified multiple open challenges that must be addressed to ensure these systems are safe, reliable, and usable in real-world settings. These include limitations in context handling, the need for persistent and structured memory, alignment with user intent, human-AI collaboration, and verification of agent behavior. As software developers are increasingly relying on AI coding agents, these concerns will become more pressing and will require interdisciplinary solutions drawing from programming languages, human-computer interaction, software engineering, and responsible AI.
Looking ahead, AI agentic programming offers exciting opportunities to fundamentally rethink the software development practice. Whether as collaborative partners in interactive workflows or as autonomous systems that manage long-running tasks, these agents have the potential to augment developer productivity, reduce software maintenance costs, and expand access to programming. We hope this survey serves as a foundation for researchers and practitioners to navigate the emerging landscape of AI agentic programming and to accelerate progress in building the next generation of intelligent and trustworthy software development tools.

### v2

@@
 Abstract.
-AI agentic programming is an emerging paradigm in which large language models (LLMs) autonomously plan, execute, and interact with external tools like compilers, debuggers, and version control systems to iteratively perform complex software development tasks.
-Unlike conventional code generation tools, agentic systems are capable of decomposing high-level goals, coordinating multi-step processes, and adapting their behavior based on intermediate feedback.
-These capabilities are transforming the software development practice.
-As this emerging field evolves rapidly, there is a need to define its scope, consolidate its technical foundations, and identify open research challenges.
-This survey provides a comprehensive and timely review of AI agentic programming.
-We introduce a taxonomy of agent behaviors and system architectures, and examine core techniques including planning, memory and context management, tool integration, and execution monitoring.
-We also analyze existing benchmarks and evaluation methodologies used to assess coding agent performance.
-Our study identifies several key challenges, including limitations in handling long context, a lack of persistent memory across tasks, and concerns around safety, alignment with user intent, and collaboration with human developers.
-We discuss emerging opportunities to improve the reliability, adaptability, and transparency of agentic systems.
-By synthesizing recent advances and outlining future directions, this survey aims to provide a foundation for research and development in building the next generation of intelligent and trustworthy AI coding agents.
+AI agentic programming is an emerging paradigm where large language model (LLM)-based coding agents autonomously plan, execute, and interact with tools such as compilers, debuggers, and version control systems.
+Unlike conventional code generation, these agents decompose goals, coordinate multi-step processes, and adapt based on feedback, reshaping software development practices.
+This survey provides a timely review of the field, introducing a taxonomy of agent behaviors and system architectures and examining relevant techniques for planning, context management, tool integration, execution monitoring, and benchmarking datasets.
+We highlight challenges of this fast-moving field and discuss opportunities for building reliable, transparent, and collaborative coding agents.
 Large Language Models, LLMs, AI Agents, AI Agentic Programming
@@
 1. Introduction
-The paradigm of software development is changing rapidly with the rise of large language models (LLMs) (hou2024largelanguagemodelssoftware, ).
-These models enable artificial intelligence (AI) systems that not only generate code (TOSEM25, ) but also understand task requirements, interact with development tools, and iteratively refine their outputs (FSE24_Alshahwan, ; qian2024chatdevcommunicativeagentssoftware, ).
-Recent studies suggest that software developers now use LLMs routinely to assist in daily coding tasks (hou2024largelanguagemodelssoftware, ; DBLP:conf/fose-ws/FanGHLSYZ23, ; gong2025language, ).
-Unlike traditional code generation tools (TraditionalCodeGenration, ) that respond to a single prompt with a static code snippet, emerging AI coding agents are designed to operate within dynamic software environments, performing iterative, tool-augmented tasks to achieve complex goals.
-This shift has given rise to a new programming paradigm known as AI agentic programming, where LLM-based coding agents can autonomously plan, execute, and refine software development tasks (gao2023retrieval, ; novikov2025alphaevolve, ).
-These agents go beyond code completion: they can generate entire programs or modules from natural language specifications, diagnose and fix bugs using compiler or test feedback, write and execute test cases, and refactor code for readability or performance.
-They can also invoke and interact with external tools such as compilers, debuggers, performance profilers, or version control systems, supporting an end-to-end development workflow.
+The software development paradigm is changing rapidly with the rise of large language models (LLMs) (Hou et al., 2024).
+These models enable artificial intelligence (AI) systems that not only translate natural language descriptions into code snippets (Jiang et al., 2025) but also understand task requirements, interact with development tools, and iteratively refine their outputs to produce complex software (Alshahwan et al., 2024; Qian et al., 2024).
+Recent studies suggest that developers now use LLMs routinely to assist in daily coding tasks (Hou et al., 2024; Fan et al., 2023b; Gong et al., 2025).
+Unlike traditional code generation tools (Allamanis et al., 2018a) that respond to a single prompt with a static code snippet, emerging AI coding agents are designed to operate within dynamic software environments, performing iterative, tool-augmented tasks to achieve complex goals.
+This shift has given rise to a new programming paradigm, AI agentic programming, where LLM-based coding agents can autonomously plan, execute, and refine software development tasks (Gao et al., 2023; Novikov et al., 2025).
+Unlike conventional code-completion tools (Int, 2025; LSP, 2025; Bruch et al., 2009a), which primarily assist with local suggestions, these agents are capable of generating entire programs or modules from natural language specifications, diagnosing and fixing bugs using compiler or test feedback, writing and executing test cases, and refactoring code for readability or performance.
+Beyond code generation, they can also orchestrate external tools, such as compilers, debuggers, performance profilers, and version control systems, supporting an end-to-end development workflow.
 This emerging paradigm has the potential to fundamentally change how software is built and maintained.
-For example, an AI agent can take a natural language description of a feature and work through a series of steps, such as writing code, generating tests, running those tests, analyzing and fixing issues, and preparing a pull request.
-Some state-of-the-art coding agents have demonstrated the ability to continue working for hours while maintaining task consistency, avoiding deadlocks, and recovering from failed actions (qian2024chatdevcommunicativeagentssoftware, ; novikov2025alphaevolve, ).
-These systems can generate and test code, migrate software between frameworks, debug runtime failures, and integrate new features by decomposing complex goals into manageable subtasks (deng2024pentestgptllmempoweredautomaticpenetration, ; liu2023bolaa, ).
+For example, an AI agent can take a natural language description of a feature and work through a series of steps, such as writing code, generating tests, running those tests, analyzing and fixing issues, and preparing a pull request (Qian et al., 2024).
+Some state-of-the-art coding agents, like Anthropic’s Calude Opus 4 (Cal, 2025), have demonstrated the ability to continue working for hours while maintaining task consistency, avoiding deadlocks, and recovering from failed actions (Qian et al., 2024; Novikov et al., 2025).
+These systems can generate and test code, migrate software between frameworks, debug runtime failures, and integrate new features by decomposing complex goals into manageable subtasks (Deng et al., 2024; Liu et al., 2023d).
 This represents a clear shift from static, one-shot AI-based code generation to interactive, iterative, and tool-augmented workflows.
@@
 There is no standard taxonomy, benchmark suite, or evaluation methodology.
-Moreover, multiple key challenges remain, including improving reliability (hou2024largelanguagemodelssoftware, ), reducing errors or hallucinations (gong2025language, ), handling tasks across different platforms and languages (DBLP:conf/pldi/polycoder22, ), and making sure these systems are safe and trustworthy in real use (yang2024swe, ).
-The success of AI coding agents depends heavily on their ability to interact effectively with external tools.
-However, today’s programming languages, compilers, and debuggers are fundamentally human-centric.
+At the same time, multiple key challenges remain.
+These include ensuring reliability and robustness in dynamic environments (Hou et al., 2024), mitigating errors and hallucinations in generated code (Gong et al., 2025), extending support beyond dominant languages such as Python to diverse platforms and software ecosystems (Xu et al., 2022), and embedding safety, trust, and accountability into autonomous behaviours (Yang et al., 2024b).
+The success of AI coding agents also depends heavily on their ability to interact effectively with external tools.
+However, today’s programming languages, compilers, and debuggers are fundamentally human-centric (Marron, 2023; Aho et al., 2006).
 They are not designed for automated, autonomous systems.
-These tools often abstract away internal states and decision-making processes to improve usability, ensure portability, and reduce cognitive load for human users (llvm, ; Maleki, ).
-While this abstraction benefits human developers, it may not fit AI agents, which require fine-grained, structured access to internal states, transformation sequences, and validation logic in order to reason about the effects of their actions (bi2024iterative, ).
+These tools often abstract away internal states and decision-making processes to improve usability, ensure portability, and reduce cognitive load for human users (Lattner and Adve, 2004a; Maleki et al., 2011b).
+While this abstraction benefits human developers, it may not fit AI agents, which require fine-grained, structured access to internal states, transformation sequences, and validation logic in order to reason about the effects of their actions (Bi et al., 2024).
 Without such access, AI agents struggle to diagnose failures, understand the implications of their changes, or recover from errors in a principled way.
-For instance, when a code transformation leads to a build failure, the agent needs more than just an error message - it must trace the failure to specific intermediate steps and understand why the changes caused the issue.
+For instance, when a code transformation leads to a build failure, the agent needs more than just an error message - it must trace the failure to specific intermediate steps and understand why certain code edits or actions caused the issue.
 Existing development environments do not provide hooks and feedback mechanisms to support this kind of iterative, tool-integrated reasoning.
-Similarly, agentic coding systems benefit significantly from toolchains that support iterative development, state tracking, and rich feedback propagation - capabilities that most conventional tools do not expose.
+Similarly, agentic coding systems would benefit from toolchains that support iterative development, state tracking, and rich feedback propagation - capabilities that most conventional tools do not expose.
 To operate effectively, AI agents may need access to internal compiler representations, transformation traces, symbolic information, and execution metadata.
-This raises a fundamental question: Are our current programming languages and software development tools still adequate in the era of AI agentic programming?
-Or is it time to rethink the design of programming languages, compilers, and debuggers to treat AI agents as first-class participants in the development process?
 These challenges show that AI agentic programming is not just a new way of using existing tools.
-It is a shift that exposes important gaps in how today’s software systems are designed.
-As the field evolves rapidly, there is a growing need to clarify its conceptual landscape, identify common patterns and system architectures, and assess the suitability of current development ecosystems.
+It is a shift that exposes important gaps in how today’s systems software is designed.
+As the field evolves rapidly, there is an urgent need to clarify its conceptual landscape, identify common patterns and system architectures, and assess the suitability of current development ecosystems.
 This is the right moment to step back, take stock of recent progress, and lay out the key questions that researchers and developers need to tackle next.
-Therefore, this survey aims to provide a comprehensive overview of the emerging field of AI agentic programming. Specifically, it covers:
+Therefore, this survey aims to provide a comprehensive overview of the emerging field of AI agentic programming.
+Specifically, it covers:
   • A conceptual foundation and taxonomy of AI coding agents,
   • A review of core system architectures and underlying techniques,
-  • A summary of current applications and practical use cases,
-  • An analysis of evaluation strategies and benchmarking practices,
+  • A summary of the behavior dimensions, operating modes, evaluation strategies and benchmarking practices of AI coding agents,
   • A discussion of key challenges and current limitations, and
   • An exploration of future research directions, including opportunities to bridge perspectives across disciplines such as programming languages, software engineering, AI, and human-computer interaction.
-We focus primarily on LLM-driven agentic systems for software development, though many insights extend to general task-oriented agents.
+We focus primarily on LLM-driven agentic systems for software development, though many insights extend to general AI agents in other domains like information retrieval (Li et al., 2025a).
 Our goal is to chart the current landscape, clarify foundational concepts, and support the design of robust, efficient, and trustworthy AI agents for programming.
@@
 2.1. AI Agentic Programming
-Figure 1. A representative workflow of an AI coding agent.
+Figure 1. An example workflow of an AI coding agent.
 AI agentic programming refers to a new programming paradigm in which LLM-based agents autonomously perform software development tasks.
-Unlike traditional code generation tools that produce outputs in a single step based on a static prompt (DBLP:conf/pldi/polycoder22, ), agentic systems operate in a goal-directed, multi-step manner.
-They reason about tasks, make decisions, use external tools (such as compilers, debuggers, and test runners), and iteratively refine their outputs based on feedback (DBLP:conf/nips/ShinnCGNY23, ; liu2024large, ; novikov2025alphaevolve, ).
+Unlike traditional code generation tools that produce outputs in a single step based on a static prompt (Xu et al., 2022), agentic systems operate in a goal-directed, multi-step manner.
+They reason about tasks, make decisions, use external tools (such as compilers, debuggers, and test runners), and iteratively refine their outputs based on feedback (Shinn et al., 2023; Liu et al., 2024e; Novikov et al., 2025).
 These agents can plan sequences of actions, adapt their strategies over time, and coordinate complex development workflows with limited or no human intervention.
-At its core, agentic programming combines the capabilities from natural language processing, external tool integration, and task planning.
-Figure 1 illustrates the architecture of a Copilot-style agentic programming system.
+At its core, agentic programming combines the capabilities of natural language processing, external tool integration, and task planning.
+Figure 1 illustrates the architecture of a GitHub Copilot-style agentic programming system (GitHub, 2021).
 At its core, the agent embeds an LLM within an execution loop, enabling interaction with the development environment.
-The agent receives natural language prompts from the user and gathers additional context from the operating system and the workspace (e.g., file summaries or environment state).
-This information is passed into the reasoning loop, where the LLM decomposes the task into subgoals, generates code or decisions, and determines whether to invoke external tools—such as reading/editing files or executing terminal commands.
+The LLM receives natural language prompts from the user and gathers additional context from the operating system and the workspace (e.g., file summaries or environment state).
+This information is passed into the reasoning loop, where the LLM decomposes the task into subgoals, generates code or decisions, and determines whether to invoke external tools like reading/editing files or executing terminal commands.
 Tool outputs are returned to the loop and used as feedback for further refinement.
@@
 Final results are streamed back to the user.
-AI agentic programming is characterized by several key properties.
+A typical AI agentic programming system is characterized by several key properties.
 First, it emphasizes autonomy, where LLM-based agents can make decisions and take actions without continuous human supervision.
@@
 Third, it supports iterative refinement, allowing agents to improve outputs based on intermediate feedback.
-Finally, it is goal-oriented, with agents pursuing high-level objectives (e.g., sub-tasks generated from the user inputs) rather than simply responding to one-shot prompts.
-Together, these features mark a departure from earlier forms of automation and code generation based on rules (xu2008rule, ), classical machine learning models (TraditionalCodeGenration, ) or one-shot LLM calling (DBLP:conf/pldi/polycoder22, ).
+Importantly, it is goal-oriented, with agents pursuing high-level objectives (e.g., sub-tasks generated from the user inputs) rather than simply responding to one-shot prompts.
+Together, these features mark a departure from earlier forms of automation and code generation based on rules (Xu, 2008), classical machine learning models (Allamanis et al., 2018a) or one-shot LLM calling (Xu et al., 2022).
 AI agentic programming represents a change toward intelligent systems that actively participate in the software development process.
-This enables new capabilities in intelligent code assistance (github_copilot, ), autonomous debugging and testing (englhardt2024exploring, ; deng2024pentestgptllmempoweredautomaticpenetration, ), automated code maintenance (fan2023automated, ), and potentially even self-improving software systems (DBLP:journals/corr/abs-2405-15189, ; DBLP:journals/corr/abs-2310-02304, ).
+This enables new capabilities in intelligent code assistance (GitHub, 2021), autonomous debugging and testing (Englhardt et al., 2024; Deng et al., 2024), automated code maintenance (Fan et al., 2023a), and potentially even self-improving software systems (Huang et al., 2024b; Zelikman et al., 2024).
 2.1.1. Working example
@@
 2.2. Historical Context and Motivations
-Figure 3. A diagram showing the evolution of coding agents from program synthesis, to code completion tools to AI coding agents
-As illustrated in Figure 3, the idea of automating software development has long been a goal in artificial intelligence and software engineering research (solar2008program, ).
-Early efforts in program synthesis aimed to generate correct-by-construction programs from formal specifications (gulwani2017program, ), while code completion tools sought to improve developer productivity by predicting likely code snippets based on contexts (github_copilot, ; tabnine, ; amazon-q-developer, ).
-These approaches, while impactful, typically relied on classical machine-learning models (TraditionalCodeGenration, ), handcrafted ruless (xu2008rule, ), or statistical techniques with limited generalizations (bruch2009learning, ).
-The advent of large-scale pre-trained LLMs such as Codexs (openai2021codex, ), GPTs (openai2024gpt4o, ), Claudes (claude2024, ), LLaMas (meta2024llama3, ) and StarCoder (lozhkov2024starcoder, ) marked a major turning point.
-These models demonstrated strong zero-shot and few-shot capabilities in generating code (DBLP:conf/nips/Le0GSH22, ; peng2024perfcodegen, ), translating between programming languages (DBLP:conf/emnlp/2020pymt5, ; DBLP:conf/mipro/GuzziKPGB24, ), and answering complex programming questions with little or no task-specific fine-tuning (DBLP:journals/corr/abs-2203-07814, ; coignion2024performance, ).
-Their ability to understand and generate natural language and code made them a good fit for software development tasks beyond basic code completion (husein2025large, ), including documentation generation (dvivedi2024comparative, ), test synthesis (jorgensen2024large, ), and bug detection and fixing (englhardt2024exploring, ; deng2024pentestgptllmempoweredautomaticpenetration, ; kang2025explainable, ).
+Figure 3. The evolution of coding agents from program synthesis, to code completion tools, to AI coding agents.
+Figure 4. Evolution of code generation approaches.
+2.2.1. A systems perspective
+As illustrated in Figure 3, the idea of automating software development has long been a goal in artificial intelligence and software engineering research (Solar-Lezama, 2008).
+In some respects, the rise of AI agents echoes the historical evolution of operating systems.
+Early operating systems introduced processes as the basic active execution units, capable of running independently.
+These evolved into threads, lightweight abstractions that enabled efficient concurrency and finer-grained scheduling (Wollheim, 1984).
+Later, with object-oriented design (Riel, 1996), objects became the dominant unit of modularity - encapsulating state and behaviour, but remaining fundamentally passive.
+Active objects (Nierstrasz, 1987, 1993) extended the object model by incorporating their own thread of control and asynchronous message passing, allowing them to act autonomously rather than merely responding to external calls—an early precursor to modern agents.
+LLM-based AI agents can be seen as the next step in this trajectory.
+Like processes and threads, they are active entities, but unlike objects, they do not wait passively for instructions at every step.
+Instead, they can plan, act, and coordinate tasks proactively across diverse contexts.
+Their behavior is not restricted to fixed, handwritten policies but is driven by LLMs, enabling them to adapt dynamically, invoke external tools, and collaborate with both humans and other agents (Li et al., 2024e).
+In this sense, agentic programming represents a new stage in the long-standing pursuit of making software development more autonomous, where code generation, debugging, and optimization are no longer fully hand-crafted but emerge through iterative interactions between humans, tools, and intelligent agents.
+2.2.2. Algorithm evolution of agents
+Figure 4 shows the evolution of algorithms.
+Early efforts in program synthesis aimed to generate correct-by-construction programs from formal specifications (Gulwani et al., 2017), while code completion tools sought to improve developer productivity by predicting likely code snippets based on contexts (GitHub, 2021; Tabnine, 2025; Amazon Web Services, 2025).
+These approaches, while impactful, typically relied on classical machine-learning models (Allamanis et al., 2018a), handcrafted ruless (Xu, 2008), or statistical techniques with limited generalizations (Bruch et al., 2009b).
+The advent of large-scale pre-trained LLMs such as Codexs (OpenAI, 2021), StarCoder (Lozhkov et al., 2024), LLaMas (Meta, 2024), GPTs (OpenAI, 2024b), Claudes (Anthropic, 2024b), Gemini (et al., 2025), Qwen (Hui et al., 2024), and Deepseek (Guo et al., 2025), marked a major turning point.
+These models demonstrated strong zero-shot and few-shot capabilities in generating code (Le et al., 2022; Peng et al., 2024a), translating between programming languages (Clement et al., 2020; Guzzi et al., 2024), and answering complex programming questions with little or no task-specific fine-tuning (Li et al., 2022; Coignion et al., 2024).
+Their ability to understand and generate natural language and code made them a good fit for software development tasks beyond basic code completion (Husein et al., 2025), including documentation generation (Dvivedi et al., 2024), test synthesis (Jorgensen et al., 2024), and bug detection and fixing (Englhardt et al., 2024; Deng et al., 2024; Kang et al., 2025).
 As these models became more capable, a new opportunity emerged: using LLMs not just as passive code generators, but as autonomous agents that could reason about goals, invoke tools, and refine their outputs over multiple steps.
@@
 Several trends motivated this change.
-First, real-world software development often requires iterative problem solving, tool use, and adaptation, which single-step code generation cannot handle effectively (liu2024large, ; li2024survey, ).
-Second, the rise of prompt engineering and structured prompting techniques (e.g., ReAct, chain-of-thought, scratchpads) enabled LLMs to reason more effectively over multiple steps (yao2023react, ; guo2025deepseek, ).
-Third, the increasing availability of APIs, command-line tools, and language server protocols made it possible to integrate LLMs into full-stack development environments (anthropic_claude_cli, ; google_cloud_cli, ; anthropic_mcp_spec, ).
-These developments prompted a rethinking of how LLMs could be deployed-not just as smart coding assistants (github_copilot, ; tabnine, ), but as semi-autonomous agents capable of carrying out software engineering tasks with minimal human supervision.
-The resulting systems promise to augment developer productivity, lower the barrier to entry for software creation, and open new avenues for automation in programming workflows.
+First, real-world software development often requires iterative problem solving, tool use, and adaptation, which single-step code generation cannot handle effectively (Liu et al., 2024e; Li et al., 2024e).
+Second, the rise of prompt engineering and structured prompting techniques (e.g., ReAct, chain-of-thought, scratchpads) enabled LLMs to reason more effectively over multiple steps (Yao et al., 2023c; Guo et al., 2025).
+Third, the increasing availability of APIs, command-line tools, and language server protocols made it possible to integrate LLMs into full-stack development environments (Anthropic, 2025a; Cloud, 2025b; Anthropic, 2025c).
+These developments prompted a rethinking of how LLMs could be deployed-not just as smart coding assistants (GitHub, 2021; Tabnine, 2025), but as semi-autonomous agents capable of carrying out software engineering tasks with minimal human supervision.
 2.3. Agency in AI Systems
@@
 In the context of AI coding agents, agency refers to a system’s capacity to act autonomously, i.e., selecting actions based on internal objectives, external feedback, and learned knowledge.
-Classical AI research has explored agency extensively in domains such as planning, robotics, and multi-agent systems (li2024survey, ; wu2024autogen, ).
-Traditional agent models focus on several key attributes:
-  • Reactivity: the ability to respond to changes in the environment,
-  • Proactivity: the pursuit of long-term goals,
-  • Social ability: the capacity to communicate and coordinate with other agents or humans,
-  • Autonomy: the ability to operate without direct human intervention.
-In the context of AI agentic programming, agency manifests through an LLM-based system’s ability to:
-  • Interpret open-ended tasks described in natural language,
-  • Plan sequences of development steps (e.g., writing, testing, debugging),
-  • Invoke and coordinate external tools (e.g., compilers, test runners, linters),
-  • Adapt actions based on feedback from the environment (e.g., compiler errors or test failures),
-  • Maintain coherent state and reasoning across multiple iterations.
+Classical AI research has explored agency extensively in domains such as planning, robotics, and multi-agent systems (Li et al., 2024e; Wu et al., 2024a).
+Traditional agent models are typically characterized by four key attributes: reactivity, the ability to respond to changes in the environment; proactivity, the pursuit of long-term goals; social ability, the capacity to communicate and coordinate with other agents or humans; and autonomy, the ability to operate without direct human intervention.
+In the context of AI agentic programming, these notions of agency are realized in new ways.
+An LLM-based system can interpret open-ended tasks expressed in natural language, plan sequences of development steps such as writing, testing, and debugging, and invoke or coordinate external tools like compilers, test runners, and linters.
+It can also adapt its actions in response to environmental feedback, such as compiler errors or test failures, while maintaining coherent state and reasoning across multiple iterations.
 Unlike symbolic AI agents that rely on explicitly defined world models and search-based planning, LLM-based coding agents operate in a probabilistic, language-driven manner.
 Despite this difference, they increasingly exhibit behaviors aligned with classical definitions of agency, especially when augmented with memory, tool-use modules, and planning routines.
-2.3.1. Comparison with Robotics and Reinforcement Learning Agents
-Robotic agents typically interact with the physical world through sensors and actuators.
-Their perception, control, and planning modules are tightly coupled with real-time feedback and often require safety guarantees.
-Reinforcement learning (RL) agents (sun2024llm, ), by contrast, learn behaviors by maximizing cumulative reward through trial and error, often in simulated environments.
-These agents explore large state-action spaces and learn policies over time.
-Agentic programming systems share similarities with both paradigms:
-  • Like robotic agents, coding agents must coordinate perception (e.g., task understanding) and action (e.g., code edits, tool invocations) in an environment with constraints and feedback.
-  • Like RL agents, agentic systems benefit from feedback loops (e.g., test results, compiler outputs) and may incorporate exploration, retry strategies, or even reward-guided behavior.
-However, coding agents operate in a symbolic, tool-rich environment, where actions are language-based, and environments (e.g., codebases, APIs, test harnesses) are highly structured.
-They must reason not only about immediate feedback but also about abstract software goals, dependencies, and long-term coherence across multiple steps.
-This makes agency in programming both uniquely challenging and distinctly different from physical-world or simulated agents.
 2.4. Key Enablers of Agentic Programming
@@
 Table 1. Representative LLMs for coding tasks.
-Model Size (B) Context Win.
-Tool use Provider (access) Open Source MoE Used in coding IDEs GPT-5 N/A 1 M OpenAI (API only) VS Code, Cursor, other IDEs GPT-4 variants (o3, o4, etc.) N/A 128k OpenAI (API only) VS Code, JetBrains, Cursor Claude 4 Opus $\sim$300 200k Anthropic (API) Cursor, Replit (chat) Gemini 2.5 Pro $\sim$200 1M Google (API) Replit, Google Colab Grok 4 $\sim$1.7T $\sim$128k xAI Not publicly integrated DeepSeek R1-0528 671 (act.
-37) 160k DeepSeek (API + weights) Emacs, VS Code (via extension) Kimi K2 1000 (act.
-32) 128k Limited Moonshot AI (API) Custom plugin support Qwen3-235B-A22B 235 (act.
-22) 32k Limited Alibaba Alibaba Cloud IDE Qwen3-Coder-480B-A35B-Instruct 480 (act.
-35) 256k Alibaba Alibaba Cloud IDE Solar-Pro 72 128k Upstage (weights on HF) VS Code (via third-party) Openhands-LM-32B-v0.1 32 128k OpenHands VS Code (via extension) Devstral-Large 123 128k Mistral (API only) VS Code (via API) Devstral-Small 24 128k Mistral (API only) VS Code (via API)
+[…]
 LLMs trained on massive corpora of code and natural language form the foundation of modern agentic programming systems.
-These models, as represented by GPT-5 (openai2025gpt5introducing, ), Claude (claude2024, ), DeepSeek (guo2025deepseek, ), and Gemini (gemini2023, ), serve as the core reasoning engines, powering code generation, task planning, debugging, documentation, and natural language interaction.
+These models, as represented by GPT-5 (OpenAI, 2025), Claude (Anthropic, 2024b), DeepSeek (Guo et al., 2025), and Gemini (Anil et al., 2023a), serve as the core reasoning engines, powering code generation, task planning, debugging, documentation, and natural language interaction.
 Their ability to understand and execute complex instructions makes them central to the design of agentic workflows.
@@
 This flexibility enables developers to use the same underlying model for a wide range of software engineering tasks, from scaffolding and unit test generation to bug repair and performance tuning.
-In addition to general-purpose models, some LLMs like Grok (grok4, ) and Calude Opus (anthropic_opus4, ) are increasingly optimized for coding tasks through specialized instruction tuning, extended context length, tool use capabilities, and integration with retrieval-based systems.
+In addition to general-purpose models, some LLMs like Grok (xAI, 2025) and Calude Opus (Anthropic, 2025b) are increasingly optimized for coding tasks through specialized instruction tuning, extended context length, tool use capabilities, and integration with retrieval-based systems.
 These enhancements make them suitable for multi-turn reasoning, code synthesis grounded in external context, and tool-augmented workflows.
-Table 1 provides a comparative overview of state-of-the-art LLMs widely used in code-related tasks.
-The models listed include Grok 4, GPT-5 (GPT-4, o3 and o4 variants), Gemini 2.5 Pro, Claude 4 Opus, DeepSeek V3, Kimi K2, and Qwen3.
-The table compares their key attributes, such as the context window length.
-As the capabilities of LLMs continue to evolve, selecting and fine-tuning the appropriate foundation model becomes a critical design choice in building reliable, efficient, and adaptive agentic systems.
+Table 1 provides a comparative overview of some of the state-of-the-art LLMs used in code-related tasks.
+The table compares their key attributes.
+As the capabilities of LLMs continue to evolve, selecting and fine-tuning the appropriate foundation model for coding tasks becomes a critical design choice in building reliable, efficient, and adaptive agentic systems.
 2.4.2. Prompt engineering and reasoning strategies
 Effective agentic behavior often requires structured prompting techniques to guide LLMs through multi-step reasoning and tool use.
-Methods such as chain-of-thought prompting, ReAct (reasoning and acting), scratchpad prompting, and modular prompting enable LLMs to plan, reflect, and revise outputs.
-These techniques allow agents to decompose complex problems, retain intermediate states, and act in ways that are more transparent and controllable.
+Rather than relying on a single input-output exchange, these methods provide scaffolding that helps models break tasks into manageable steps and maintain coherence across longer interactions.
+For instance, chain of thought prompting (Wei et al., 2022) encourages explicit reasoning traces, making intermediate steps visible and improving problem-solving accuracy.
+ReAct (reasoning and acting) interleaves reasoning with concrete actions (Yao et al., 2023c), such as tool calls or environment interactions, enabling agents to both deliberate and act in context.
+Scratchpad prompting (Abbe et al., 2024) provides a working memory space where partial results, hypotheses, or plans can be written down and refined, supporting iterative refinement.
+Modular prompting (Sun et al., 2023a), on the other hand, separates tasks into distinct functional roles—such as planner, executor, and verifier—so that the model can coordinate across specialized subtasks.
+Together, these techniques allow agents to decompose complex problems, retain intermediate states, and revise their behavior in light of new evidence.
+They also increase transparency by exposing the reasoning process and provide greater controllability by constraining how models structure their outputs.
+In practice, structured prompting forms the backbone of many agentic systems, enabling LLMs to move beyond ad hoc responses and toward more reliable, interpretable, and goal-directed behavior.
 2.4.3. Tool use and API integration
 Table 2. Examples of tools supported by GitHub Copilot agent.
-Tool Type Examples Compiler gcc (GCC, ), clang (Clang, ), javac (Javac, ), tsc (TSC, ) Debugger gdb (GDB, ), lldb (LLDB, ), pdb (PDB, ) Test Framework pytest (Pytest, ), unittest (Unittest, ), Jest (Jest, ), Mocha (Mocha, ) Linter eslint (ESLint, ), flake8 (Flake8, ), black (Black, ), prettier (Prettier, ) Version Control git (Git, ) Build System make (Make, ), cmake (CMake, ), npm (NPM, ), maven (Maven, ) Package Manager pip (Pip, ) yarn (Yarn, ), cargo (Cargo, ) Language Server pyright (Pyright, ), tsserver (TSServer, )
+[…]
+import OpenAI from "openai";
+const client = new OpenAI();
+const tools = [
+"type": "function",
+"function": {
+"name": "compile_code",
+"description": "...",
+"parameters": {
+"type": "object",
+"properties": {
+"language": {
+"type": "string",
+"enum": ["c", "cpp"],
+"description": "Programming language to compile."
+"source": {
+"type": "string",
+"description": "Source code to compile."
+"flags": {
+"type": "array",
+"items": { "type": "string" },
+"description": "Compiler flags."
+"required": ["language", "source"],
+"additionalProperties": false
+Listing 1: An example of the OpenAI tool schema
+Table 3. Example interfaces between LLMs and tools.
+[…]
 Agentic systems rely heavily on external tools, such as compilers, debuggers, test frameworks, linters, and version control systems, to validate and refine generated code.
-Integration with these tools via command-line interfaces, language server protocols (LSP), or RESTful APIs enables agents to execute actions, gather feedback, and close the loop between generation and validation.
-Tool use is a defining feature of agentic programming, supporting iterative workflows and grounding decisions in observable outcomes.
-For example, Table 2 lists some of the tools supported by GitHub Copilot Agent. The set of tools ranges from compilation to testing and version control.
+These tools provide the concrete signals needed to check correctness, enforce coding standards, and ensure that outputs remain consistent with project requirements.
+For example, Table 2 lists a subset of the tools currently supported by the GitHub Copilot Agent (GitHub, 2021), covering compilation, testing, and version control.
+Integration with external tools can take multiple forms, including command line interfaces, language server protocols (LSP), and RESTful APIs.
+Increasingly, LLMs interact with tools through structured Python or JavaScript interfaces that specify the available actions, input parameters, and expected outputs in a machine-readable format.
+For example, Listing 1 shows how a compiler can be exposed as a tool to extend an LLM’s capabilities.
+This structured approach reduces ambiguity, grounds commands in the correct syntax, and makes it easier for the model to call external tools safely and consistently.
+By interpreting the schema, an LLM can generate well-formed commands, parse structured responses, and adjust its behavior in a predictable way.
+Table 3 summarizes the main types of interfaces through which LLMs interact with external tools.
+These range from free-form natural language instructions to highly structured schemas and domain-specific protocols.
 2.4.4. State and context management
@@
 Table 4 compares the context management strategies of mainstream AI coding agents, revealing substantial differences in context size and memory persistence.
-Tools like GitHub Copilot and Codeium lack persistent memory, instead using transient methods such as sliding windows or dynamic token budgeting.
-In contrast, agents like SWE-agent, Devika, and OpenDevin employ persistent storage, often via vector databases or structured stores, to support long-term recall of plans, tool outputs, and project history.
-Some, such as Cursor IDE and Continue.dev, use embedding-based search to retrieve semantically relevant content, while others summarize prior actions to stay within the available context window.
+Tools like GitHub Copilot (GitHub, 2021) currently do not utilize persistent memory, instead using transient methods such as sliding windows or dynamic token budgeting.
+In contrast, agents like SWE-agent (Yang et al., 2024b), Devika (dev, 2025), and OpenDevin (Ope, 2025) employ persistent storage, often via vector databases or structured stores, to support long-term recall of plans, tool outputs, and project history.
+Some, such as Cursor IDE (Cursor, 2023) and Continue.dev (con, 2025), use embedding-based search to retrieve semantically relevant content, while others summarize prior actions to stay within the available context window.
 These differences reflect a clear trade-off: smaller context windows typically rely on lightweight retrieval or summarization, whereas larger windows with persistent memory enable richer state tracking but add storage and retrieval overhead.
@@
 Agents may rerun failed tests, revise prompts based on compiler errors, or reflect on past failures to improve future behavior.
-Some systems incorporate explicit planning, retry mechanisms, or even gradient-based updates through fine-tuning or reinforcement learning.
+For instance, compiler errors may trigger targeted code edits, test failures may prompt iterative debugging, and linter warnings may guide stylistic refinements.
+Some systems incorporate explicit planning, retry mechanisms, or even gradient-based updates (Huang et al., 2024c) through fine-tuning or reinforcement learning.
 This closed-loop design supports robustness and adaptability in complex programming tasks.
-Together, these enablers mark a significant departure from traditional code generation pipelines.
-They allow LLMs to operate as interactive systems capable of planning, acting, and learning, paving the way for more autonomous and capable software development agents.
 2.5. Comparison to Related Paradigms
@@
 2.5.1. Program synthesis
-Program synthesis has been a foundational approach to automated code generation, traditionally divided into two types: deductive synthesis uses formal specifications to generate provably correct programs, while inductive synthesis learns from input-output examples with symbolic search and logic programming techniques to infer program logic (gulwani2017program, ; jain2022jigsaw, ).
-Classical synthesis systems like sketching (solar2008program, ) and more recent neural approaches like RobustFill (devlin2017robustfill, ) specialize in generating targeted code snippets that satisfy precise specifications.
-However, program synthesis focuses on single-function generation from formal specifications and typically operates in a one-shot generation mode (devlin2017robustfill, ), whereas agentic programming handles multi-step workflows (e.g., planning, tool use, and iterative refinement) and engages in continuous interaction with development environments (bouzenia2024repairagent, ).
+Program synthesis has been a foundational approach to automated code generation, traditionally divided into two types: deductive synthesis uses formal specifications to generate provably correct programs, while inductive synthesis learns from input-output examples with symbolic search and logic programming techniques to infer program logic (Gulwani et al., 2017; Jain et al., 2022).
+Classical synthesis systems like sketching (Solar-Lezama, 2008) and more recent neural approaches like RobustFill (Devlin et al., 2017) specialize in generating targeted code snippets that satisfy precise specifications.
+However, classical program synthesis focuses on single-function generation from formal specifications (due to the scaling challenge of code sysnthesis (Gulwani et al., 2017)) and typically operates in a one-shot generation mode (Devlin et al., 2017), whereas agentic programming handles multi-step workflows (e.g., planning, tool use, and iterative refinement) and engages in continuous interaction with development environments (Bouzenia et al., 2024).
 Table 4. Context management mechanisms supported by mainstream AI coding agents.
-Agent Context Window Persistent Memory Context Management Mechanism GitHub Copilot 16k Sliding window over active buffer Codeium 32k Dynamic token budgeting based on file proximity and edit history Cursor IDE 128k Semantic search over project history SWE-agent 16k Vector DB retrieval for tool outputs and plan state Devika 32k Structured memory via SQLite and embeddings AutoDev 16k summarization of prior actions and tool logs Continue.dev 32k Embedding-based local recall over recent edits OpenDevin 32k RAG over command history, plans, and intermediate outputs
+[…]
 2.5.2. Code completion tools
-Code completion, as one of the most commercially successful applications of AI in programming, excels at context-aware code suggestion, leveraging large-scale pre-training on code repositories to predict next tokens of partially written code (DBLP:conf/iclr/codegen23, ; DBLP:journals/corr/abs-2203-07814, ; husein2025large, ).
-Advanced completion tools can suggest entire functions, classes, or small modules based on comments, function signatures, and surrounding context, with tools like GitHub Copilot (github_copilot, ), TabNine (tabnine, ), and Amazon Q Developer (amazon-q-developer, ) achieving widespread adoption.
+Code completion, as one of the most commercially successful applications of AI in programming, excels at context-aware code suggestion, leveraging large-scale pre-training on code repositories to predict next tokens of partially written code (Nijkamp et al., 2023; Li et al., 2022; Husein et al., 2025).
+Advanced completion tools can suggest entire functions, classes, or small modules based on comments, function signatures, and surrounding context, with tools like GitHub Copilot (GitHub, 2021), TabNine (Tabnine, 2025), and Amazon Q Developer (Amazon Web Services, 2025) achieving widespread adoption.
 These code completion tools often operate as reactive assistants that respond to developer input, while agentic programming systems demonstrate proactive behavior and autonomous planning.
-Furthermore, agentic programming extends beyond code generation to encompass testing, debugging, deployment, and maintenance activities that completion tools typically do not address (yang2024swe, ; bouzenia2024repairagent, ).
+Furthermore, agentic programming extends beyond code generation to encompass testing, debugging, deployment, and maintenance activities that completion tools typically do not address (Yang et al., 2024b; Bouzenia et al., 2024).
 2.5.3. DevOps automation
-DevOps automation focuses on streamlining software delivery pipelines through Infrastructure as Code (IaC), Continuous Integration/Continuous Deployment (CI/CD), and automated testing frameworks (humble2010continuous, ).
-Tools like Jenkins (jenkins, ), GitLab CI (gitlab_ci, ), and modern platforms like GitHub Actions (github_actions, ) automate repetitive deployment tasks, testing workflows, and infrastructure management.
+DevOps automation focuses on streamlining software delivery pipelines through Infrastructure as Code (IaC), Continuous Integration/Continuous Deployment (CI/CD), and automated testing frameworks (Humble and Farley, 2010).
+Tools like Jenkins (Project, 2025), GitLab CI (Inc., 2025), and modern platforms like GitHub Actions (GitHub, 2025) automate repetitive deployment tasks, testing workflows, and infrastructure management.
 While both paradigms emphasize automation, DevOps automation primarily handles pre-defined workflows and infrastructure management, whereas agentic programming focuses on adaptive problem-solving and creative solution generation.
-Additionally, agentic programming can potentially orchestrate and improve DevOps processes themselves, representing a higher-order form of automation (kambala2024intelligent, ).
-2.5.4. AutoML and Automated Development
-Automated Machine Learning (AutoML) represents a successful paradigm for democratizing AI model development through automation of model selection, hyperparameter tuning, and feature engineering (he2021automl, ).
-Platforms like Google Cloud AutoML (google_automl, ), Amazon SageMaker Autopilot (sagemaker_autopilot, ), and open-source frameworks like Auto-sklearn automate the traditional machine learning pipeline from data preprocessing to model deployment (feurer2015efficient, ).
+Additionally, agentic programming can potentially orchestrate and improve DevOps processes themselves, representing a higher-order form of automation (KAMBALA, 2024).
+2.5.4. Automated machine learning and automated development
+Automated Machine Learning (AutoML) represents a successful paradigm for democratizing AI model development through automation of model selection, hyperparameter tuning, and feature engineering (He et al., 2021).
+Platforms like Google Cloud AutoML (Cloud, 2025a), Amazon SageMaker Autopilot (Services, 2025), and open-source frameworks like Auto-sklearn automate the traditional machine learning pipeline from data preprocessing to model deployment (Feurer et al., 2015).
 However, AutoML focuses on statistical model optimization within well-defined machine learning workflows and operates with structured data and standardized evaluation metrics, while agentic programming tackles more general software development challenges with creative problem-solving and multi-modal reasoning.
 2.5.5. Multi-Agent systems and human-AI collaboration
-Traditional multi-agent systems in software development typically involve specialized agent roles working within predefined coordination protocols (jennings2000agent, ).
+Traditional multi-agent systems in software development typically involve specialized agent roles working within predefined coordination protocols (Jennings, 2000).
 These systems often feature separate agents for requirements analysis, code generation, testing, and documentation, coordinating through structured communication interfaces.
-Recent advances in LLM-based multi-agent programming have demonstrated the potential for more sophisticated collaboration, with systems like MetaGPT showing how multiple AI agents can simulate software development teams (hong2023metagpt, ).
+Recent advances in LLM-based multi-agent programming have demonstrated the potential for more sophisticated collaboration, with systems like MetaGPT showing how multiple AI agents can simulate software development teams (Hong et al., 2023).
 AI agentic programming can be viewed as an evolution of multi-agent systems that incorporates human-in-the-loop collaboration and dynamic role adaptation.
@@
 Moreover, the integration of tool use, environmental interaction, and persistent memory distinguishes modern agentic programming from earlier multi-agent approaches.
-Contemporary agentic systems like AutoGen (wu2024autogen, ) and CrewAI (duan2024exploration, ) enable agents to directly interact with development tools, maintain context across extended sessions, and learn from past interactions (li2024survey, ).
+Contemporary agentic systems like AutoGen (Wu et al., 2024a) and CrewAI (Duan and Wang, 2024) enable agents to directly interact with development tools, maintain context across extended sessions, and learn from past interactions (Li et al., 2024e).
 This represents a significant advancement over traditional multi-agent systems that typically operated in more constrained, simulation-based environments.
+2.5.6. Comparison with robotics and reinforcement learning agents
+Robotic agents typically interact with the physical world through sensors and actuators.
+Their perception, control, and planning modules are tightly coupled with real-time feedback and often require safety guarantees.
+Reinforcement learning (RL) agents (Sun et al., 2024c), by contrast, learn behaviors by maximizing cumulative reward through trial and error, often in simulated environments.
+These agents explore large state-action spaces and acquire policies over time.
+Agentic programming systems share features with both paradigms.
+Like robotic agents, coding agents must coordinate perception, such as task understanding, with actions such as code edits or tool invocations, all within an environment shaped by constraints and feedback.
+Like RL agents, they benefit from feedback loops, for example, test results or compiler outputs, and may employ exploration, retry strategies, or even reward-guided behavior.
+At the same time, coding agents differ in multiple ways.
+They operate in a symbolic, tool-rich environment where actions are language-based and environments, such as codebases, APIs, and test harnesses, are highly structured.
+Success requires reasoning not only about immediate feedback but also about abstract software goals, dependencies, and long-term coherence across multiple steps.
+This makes agency in programming uniquely challenging and distinct from both the physical world and simulated agents.
 3. Survey Methodology
-This survey follows a widely-used systematic literature review (SLR) methodology (keele2007guidelines, ; DBLP:journals/corr/Gong24Deep, ; hou2024largelanguagemodelssoftware, ; DBLP:journals/tse/WangHGGZFSLZN23, ; zhang2024systematic, ; qudus2025fact, ) to provide comprehensive coverage of AI agentic programming research, as illustrated in Figure 5.
+This survey follows a widely-used systematic literature review (SLR) methodology (Kitchenham and Charters, 2007; Gong and Chen, 2024a; Hou et al., 2024; Wang et al., 2023b; Zhang et al., 2024a; Qudus et al., 2025) to provide comprehensive coverage of AI agentic programming research, as illustrated in Figure 6.
+Figure 5. Survey methodology for academic research.
+[…]
+[width=]figures/pie_year
+Figure 6. Distribution of academic papers.
 3.1. Search Strategy
@@
 We also examined proceedings from top-tier venues (FSE, ICSE, ASE, ICML, NeurIPS, AAAI, etc.).
+As this fast-evolving field is largely dominated by industry, we also pay attention to open-source and the industry releases of relevant results and tools.
 Our search string combined the following term clusters using Boolean operators:
@@
 3.2. Study Selection
-After initial retrieval, we followed a three-stage study selection process: (1) title and abstract screening by two independent researchers, (2) full-text review with disagreement resolution through discussion, and (3) backward and forward citation chaining to identify additional relevant studies.
+After initial retrieval, we followed a three-stage study selection process for academic papers: (1) title and abstract screening by two independent researchers, (2) full-text review with disagreement resolution through discussion, and (3) backward and forward citation chaining to identify additional relevant studies.
 During the selection process, we used the following criteria:
@@
   • Final corpus: 152 papers included after full-text evaluation and citation chaining
+  • Software tools and industry products: we also study a wide range of state-of-the-art AI coding agents and LLMs like GitHub Copilot Agents, GPT, Gemini, Deepseek, Qwen and Claude Opus 4.
 Among the 152 academic references published from 2022 to 2025 (excluding tool descriptions and websites), 5% appeared in 2022, 22% in 2023, 53% in 2024, and 20% in 2025, as shown in Figure 6, reflecting a surge in AI agent programming research following the widespread adoption of LLMs.
-Figure 4. Survey Methodology.
+Based on our systematic analysis, we developed a hierarchical classification of AI agentic programming systems along behavioral characteristics and system architectures, as shown in Figure 7.
+The following section examines each category in detail.
 […]
-[width=]figures/pie_year
-Figure 5. Distribution of Papers.
+[width=]figures/tree
+Figure 7. Taxonomy of AI agentic programming systems.
 4. Taxonomy of AI Agentic Programming
@@
 In contrast, multi-turn agents, such as GitHub Copilot Agent or Claude Opus 4 with tooling capabilities, maintain state across interactions, enabling iterative refinement, exploration, and goal pursuit.
-For instance, GitHub Copilot Agent can hold a conversation across multiple steps, remember earlier function names, and build a complete module through back-and-forth iterations between agents (github_copilot, ).
+For instance, GitHub Copilot Agent can hold a conversation across multiple steps, remember earlier function names, and build a complete module through back-and-forth iterations between agents (GitHub, 2021).
 4.1.3. Tool-augmented vs. standalone agents
@@
 4.1.4. Static vs. adaptive agents
-Static agents (e.g., GitHub Copilot and Tabnine) follow predefined workflows or heuristics.
+Static agents (e.g., GitHub Copilot and Tabnine (Tabnine, 2025)) follow predefined workflows or heuristics.
 Adaptive agents modify their strategies using feedback from tools, user input, or environmental signals.
@@
 They are typically integrated directly into editors and IDEs, where developers interact with the underlying LLMs either through chat-like interfaces or by selecting code or comments using mouse-based interactions.
-GitHub Copilot (github_copilot, ) and Cursor (cursor, ) are two representative examples of LLM-based code assistants.
-GitHub Copilot, originally developed based on Codex trained on GitHub code repositories, offers context-aware code completions across multiple programming languages and is tightly integrated into popular IDEs such as Visual Studio Code (vscode, ) and JetBrains.
-Cursor extends this functionality by embedding conversational interaction, maintaining memory of previous edits, and supporting structured command execution, allowing for a more dynamic and iterative user experience.
-Other notable systems include Amazon Q Develiper (amazon-q-developer, ) and Tabnine (tabnine, ).
-Amazon CodeWhisperer targets developers working in cloud ecosystems, offering language-specific completions, cloud API integration, and basic vulnerability detection.
-In contrast to CodeWhisperer, Tabnine takes a privacy-first approach by deploying smaller local models trained on permissively licensed code, making it attractive for organizations and developers who do not want to send their code to a remote cloud.
+GitHub Copilot (GitHub, 2021) and Cursor (cur, 2025) are two representative examples of LLM-based code assistants.
+GitHub Copilot, originally developed based on Codex trained on GitHub code repositories, offers context-aware code completions across multiple programming languages and is tightly integrated into popular IDEs like Visual Studio Code (vsc, 2025) and JetBrains.
+Cursor extends this functionality by embedding conversational interaction, maintaining memory of previous edits, and supporting structured command execution.
+Other notable systems include Amazon Q Developer (Amazon Web Services, 2025) and Tabnine (Tabnine, 2025).
+Amazon Q Developer targets developers working in cloud ecosystems, offering language-specific completions, cloud API integration, and basic vulnerability detection.
+In contrast to Amazon Q Developer, Tabnine takes a privacy-first approach by deploying smaller local models trained on permissively licensed code, making it attractive for organizations and developers who do not want to send their code to a remote cloud.
 Implementations of these systems typically exhibit reactive behavior, responding to user input without initiating their own plans or taking proactive steps.
@@
 Many integrate external tools such as debuggers, package managers, or search engines, enabling them to gather information, resolve errors, and optimize code without direct user guidance.
-Examples include the state-of-the-art LLMs that have tool integration capabilities like GPT-5 (openai2025gpt5introducing, ), Claude Opus 4 (anthropic_opus4, ), Google Jules (Jules, ), Kimi K2 (team2025kimi, ), and Gemini 2 (comanici2025gemini25pushingfrontier, ).
+Examples include GPT-5 (OpenAI, 2025), which functions as an end-to-end collaborator for long-horizon code generation and debugging; Claude Opus 4 (Anthropic, 2025b), engineered for sustained agentic workflows with strong long-context and tool-integration capabilities; Google Jules (Jul, 2025), which can sandbox repositories, propose diffs, and execute verified changes on real projects; DevGPT (devgpt-labs, 2024), an open-source pipeline that converts tickets into actionable code and CI artifacts; Kimi K2 (Team et al., 2025), a model optimized for agentic coding proficiency; and Gemini 2 (et al., 2025), which emphasizes multimodal inputs and built-in planning modes.
 These agents are often proactive in suggesting next steps, adaptive to new inputs, and capable of maintaining continuity across extended sessions through persistent memory or context management mechanisms.
@@
 This approach improves the handling of long-horizon tasks.
-For example, CAMEL (li2023camel, ) employs two agents in a role-playing setup, a “user” agent and an “assistant” agent, which collaboratively refine goals and strategies, producing plans that downstream code generators can execute.
+For example, CAMEL (Li et al., 2023b) employs two agents in a role-playing setup, a “user” agent and an “assistant” agent, which collaboratively refine goals and strategies, producing plans that downstream code generators can execute.
+Voyager (Wang et al., 2023c) demonstrates this paradigm by continuously exploring an open-ended environment, generating executable skills, and refining them into reusable plans for long-term adaptability.
+Similarly, CodePlan (Wen et al., 2024) introduces code-form planning, where structured pseudocode serves as an explicit intermediate representation to decompose complex problems into executable steps.
 These agents are typically multi-turn, memory-enabled, and trading speed for robustness.
@@
 This approach draws inspiration from human software teams, where each member has a distinct role, such as requirements analysis, coding, testing, or documentation, and communication protocols are established to ensure progress toward shared objectives.
-For example, SWE-Agent (yang2024swe, ) employs multiple role-specific LLM agents—an “Architect” agent for high-level design, a “Coder” agent for implementation, and a “Reviewer” agent for quality assurance—connected through structured dialogue and shared memory.
-Such systems offer high autonomy, scalability, and tool integration, but face challenges in avoiding redundancy and managing conflicts.
+For example, SWE-Agent (Yang et al., 2024b) employs multiple role-specific LLM agents: an “Architect” agent for high-level design, a “Coder” agent for implementation, and a “Reviewer” agent for quality assurance, which are connected through structured dialogue and shared memory.
+ChatDev (Qian et al., 2024) follows this paradigm by simulating an end-to-end software company, where agents take on roles such as CEO, CTO, and programmers to collaboratively design, implement, and test applications.
+Furthermore, AutoCodeRover (Zhang et al., 2024e) extends collaboration into real-world repositories, orchestrating specialized agents that autonomously navigate, edit, and validate source code across complex multi-file projects.
 4.3. Summary and Comparative Table
@@
 System Category Proactivity Multi-turn Tool Use Adaptivity
-GitHub Copilot IDE Assistant Reactive
-Tabnine IDE Assistant Reactive
-Cursor IDE Assistant Reactive
-Claude Opus 4 Task-oriented Proactive
-Google Jules Task-oriented Proactive
-Dev-GPT Task-oriented Proactive
-KimI K2 Task-oriented Proactive
-Gemini 2 Task-oriented Proactive
-Voyager Planning Agent Proactive
-CAMEL Planning Agent Proactive
-CodePlan Planning Agent Proactive
-ChatDev Multi-agent System Proactive
-AutoCodeRover Multi-agent System Proactive
-SWE-Agent Multi-agent System Proactive
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 4.4. Cost and Token Consumption Model
@@
 Table 6 summarizes representative pricing.
-Table 5. Representative token pricing for LLMs used in coding tasks (USD per 1M tokens, as of Aug. 2025).
+Table 6. Example token pricing for LLMs used in coding tasks (USD per 1M tokens, as of Sep. 2025).
 Model Input(\$) Output(\$) Context Window
 GPT-5 (Standard) 1.25 10.00 256k
-GPT-5 Mini 0.25 2.00 128k
+[…]
 GPT-4 variants (e.g., 4o) 2.50 10.00 128k
-Claude 4 Opus 15.00 75.00 200k
+[…]
 Gemini 2.5 Pro 1.25 10 1M
-Grok 4 3.00 15.00 256k
+[…]
 DeepSeek R1-0528 0.55 2.19 160k
-Kimi K2 0.15 2.50 128k
-Qwen3-235B-A22B 0.22 0.88 32k
-Qwen3-Coder-480B-A35B-Instruct 4.5 13.5 256k
-Solar-Pro 0.30 0.30 128k
+[…]
+Qwen3-235B-A22B 0.22 0.88 128k
+[…]
+[…]
 Openhands-LM-32B-v0.1 2.6 3.4 128k
-Devstral-Small (Mistral) 0.10 0.30 128k
-Devstral-Large (Mistral Medium 3) 0.40 2.00 128k
+[…]
+Devstral-Medium 0.40 2.00 128k
 4.4.2. Impact of reasoning strategies.
@@
 AI agentic programming introduces a promising and complex shift in how software is developed, relying on the autonomous capabilities of LLMs.
-Despite recent progress, several technical and conceptual challenges remain that hinder the deployment of robust, scalable, and trustworthy agentic systems (sapkota2025aiagentsvsagentic, ; acharya2025agentic, ).
-5.1. Inadequate Evaluation and Benchmarking
-A variety of benchmarks and open-source toolkits (liu2023agentbench, ; qin2023toolllmfacilitatinglargelanguage, ) have been proposed to evaluate the capabilities of LLM-based agents across different tasks.
-Despite this progress, existing benchmarks commonly used for assessing coding agents, such as HumanEval (zheng2023codegeex, ) and SWE-Bench (jimenez2023swe, ), remain inadequate for capturing the full complexity of real-world software engineering workflows.
-These benchmarks primarily focus on small, self-contained problems, often restricted to a narrow range of programming languages, such as Python (chen2021evaluatinglargelanguagemodels, ), and generally lack support for interactive, multi-turn, or tool-integrated tasks (liu2023agentbench, ).
+Despite recent progress, several technical and conceptual challenges remain that hinder the deployment of robust, scalable, and trustworthy agentic systems (Sapkota et al., 2025; Acharya et al., 2025).
+5.1. Evaluation and Benchmarking
+A variety of benchmarks and open-source toolkits (Liu et al., 2023e; Qin et al., 2023) have been proposed to evaluate the capabilities of LLM-based agents across different tasks.
+Despite this progress, existing benchmarks commonly used for assessing coding agents, such as HumanEval (Zheng et al., 2023b) and SWE-Bench (Jimenez et al., 2023), may still be inadequate for capturing the full complexity of real-world software engineering workflows.
+Most of these benchmarks primarily focus on small, self-contained problems, often restricted to a small number of programming languages, such as Python (Chen et al., 2021a), and generally lack support for interactive, multi-turn, or tool-integrated tasks (Liu et al., 2023e).
 In contrast, practical agentic systems are expected to operate over large, modular codebases, interface with third-party libraries, manage build workflow pipelines, and respond dynamically to user feedback or runtime tool outputs.
-As LLM-based systems increasingly incorporate reinforcement learning (ouyang2022training, ; wang2024survey, ) and more advanced planning mechanisms, future benchmarks should reflect this integration.
-Moreover, there is a noticeable absence of evaluation frameworks designed for emerging complex use cases, such as those involving interactions with compilers and debuggers (wang2024executablecodeactionselicit, ), where agents must reason about low-level program behavior, perform iterative transformations, or track state across toolchains.
+As LLM-based systems increasingly incorporate reinforcement learning (Ouyang et al., 2022; Wang et al., 2024e) and more advanced planning mechanisms, future benchmarks should reflect this integration.
+Table 7 shows the characteristics of tasks in SWE-Bench.
+Although SWE-Bench introduces project-level repositories and leverages unit tests and continuous integration (CI) for evaluation, its scope is restricted to Python.
+Most tasks are function- or module-level, with no support for multi-turn feedback, third-party library usage, or build pipeline management.
+Even for project-level tasks, interaction is limited to binary pass/fail outcomes, providing only minimal support for realistic multi-step or tool-integrated software engineering workflows.
+Moreover, there is a noticeable absence of evaluation frameworks designed for emerging complex use cases, such as those involving interactions with compilers and debuggers (Wang et al., 2024b), where agents must reason about low-level program behavior, perform iterative transformations, or track state across toolchains.
 The lack of such domain-specific benchmarks presents a significant gap in evaluating agent performance under realistic conditions.
-5.2. Domain Foundation Models for Agents
-Generic coding agents often struggle in domain-specific environments such as embedded systems, high-performance computing, optimization, or formal software verification (lecong2025llmsreasonprogramsemantics, ; loughridge2024dafnybench, ).
+Table 7. SWE-Bench characteristics.
+Task type Proportion Languages Interaction Multi-turn feedback Library integration Build pipeline
+Function-level  65% Python Unit tests
+[…]
+Project-level ¡10% Python Tests and CI Limited (pass/fail)
+5.2. Communication Protocols for Multi-Agent Systems
+Early protocols enabled one-to-one agent–to–tool interactions but did not support direct agent-to-agent communication (MCP, 2025).
+Current practice often relies on heterogeneous web service protocols and adapters, which provide interoperability but introduce high latency, bandwidth overhead, and limited scalability (Naik and Jenkins, 2016).
+Recent work (Xu et al., 2012; A2A, 2025; Verma et al., 2015) explores group session models, where each execution unit functions as a session or group agent coordinating multiple services, and standards efforts propose unified message formats and session semantics.
+However, the lack of a common protocol, inefficiencies of adapters, and challenges in managing dynamic multi-agent sessions remain open problems for scalable and dependable agentic systems.
+5.3. Domain-specific Models for Agents
+Generic coding agents often struggle in domain-specific environments such as embedded systems, high-performance computing, optimization, or formal software verification (Le-Cong et al., 2025; Loughridge et al., 2024).
 These domains typically impose stricter operational constraints and require deep integration with specialized APIs, toolchains, and domain knowledge resources that are often underrepresented in general-purpose training corpora.
-To address these limitations, recent research has proposed domain-adapted models and task-specific learning strategies to accelerate agent performance in specialized settings (song2025injecting, ).
-For instance, some approaches have begun incorporating compiler knowledge or security-specific patterns into LLM training pipelines (lin2025llmssecuritycoursesecuring, ; LLMCompiler, ), enabling agents to reason more effectively about low-level program behavior or vulnerability patterns.
+To address these limitations, recent research has proposed domain-adapted models and task-specific learning strategies to accelerate agent performance in specialized settings (Song et al., 2025).
+For instance, some approaches have begun incorporating compiler knowledge or security-specific patterns into LLM training pipelines (Lin et al., 2025; Cummins et al., 2025), enabling agents to reason more effectively about low-level program behavior or vulnerability patterns.
 In the future, developing robust and adaptable domain foundation models will be a promising direction for enabling agents to operate reliably in complex software environments, such as LLMs pretrained or fine-tuned on domain-specific data, tools, and semantics.
@@
 As agentic systems gain increasing autonomy, so does the potential for unsafe behavior.
-Unlike traditional tools (Checkmarx, ), agentic systems can invoke external tools, perform structural code modifications, and even commit changes without direct human oversight (triedman2025multiagentsystemsexecutearbitrary, ; zhang2025agentsafetybenchevaluatingsafetyllm, ).
+Unlike traditional tools (Che, 2025), agentic systems can invoke external tools, perform structural code modifications, and even commit changes without direct human oversight (Triedman et al., 2025; Zhang et al., 2025).
 These capabilities introduce significant risks, including the possibility of introducing subtle bugs, propagating unsafe patterns, or violating security constraints.
@@
 For example, when agents visit private repositories or are deployed in cloud-integrated environments, future models may need built-in controls to restrict access to sensitive project data.
-Further, malicious prompts, poisoned APIs, or compromised toolchains can mislead agents into executing unsafe behaviors (debenedetti2024agentdojodynamicenvironmentevaluate, ; zhan2024injecagentbenchmarkingindirectprompt, ).
+Further, malicious prompts, poisoned APIs, or compromised toolchains can mislead agents into executing unsafe behaviors (Debenedetti et al., 2024; Zhan et al., 2024).
 Future research should prioritize the design of secure protocols for agent collaboration, including authentication between agents, validation of tool outputs, and detection of anomalous actions.
@@
 5.5. Toolchain Integration and Programming Language Design
-One fundamental challenge lies in the incompatibility between existing software tools and the needs of autonomous agents.
-Most programming languages, compilers, debuggers, and development tools are designed with human developers in mind (rorseth2025ladybug, ).
-They emphasize usability over feedback, so agents often struggle to diagnose failures, trace the effects of their code transformations, or understand build errors (kim2024llm, ).
-For example, compilers typically provide minimal insights into transformation failures or semantic conflicts (li2025review, ).
-To support agentic workflows, toolchains must evolve to expose IRs, transformation traces, and structured feedback interfaces.
-Moreover, language-level annotations and agent-aware interfaces are needed to communicate developer intent and guide automated reasoning.
+One fundamental challenge is the incompatibility between existing software tools and the needs of autonomous coding agents.
+Most programming languages, compilers, debuggers, and development environments were designed for human developers (Rorseth et al., 2025).
+They emphasize usability and readability over structured, machine-consumable feedback.
+As a result, agents often struggle to diagnose failures, trace the consequences of code transformations, or interpret build errors (Kim et al., 2024).
+For example, compilers typically report transformation failures or semantic conflicts in the form of coarse error messages, providing little insight into why an optimization was blocked or how a type error emerged (Li, 2025).
+Languages similarly prioritize human readability over machine-negotiated meaning, while compilers conceal internal reasoning to avoid overwhelming human users.
+These design choices, while historically effective, now limit the ability of AI coding agents to construct safe, efficient, and adaptive software.
+To enable agentic workflows, toolchains should evolve to expose richer intermediate representations, transformation traces, and structured feedback interfaces.
+Equally important, programming languages should incorporate annotations and agent-aware interfaces that make developer intent explicit, allowing automated reasoning to be guided by semantic contracts rather than inferred heuristics.
+Together, these advances would transform compilers and languages from opaque tools into collaborative infrastructures capable of supporting autonomous agents at scale.
 5.6. Scalable Memory
 Agentic programming systems must maintain coherence and reasoning over long-running tasks involving multiple iterations, tools, and contextual dependencies.
-However, current LLMs are limited by fixed context windows and lack persistent, structured memory mechanisms (wang2023augmenting, ).
-Realistic software tasks, such as feature implementation, debugging, or refactoring, require agents to store and reason over evolving states, feedback logs, intermediate plans, and prior actions (qian2025memorag, ).
+However, current LLMs are limited by fixed context windows and lack persistent, structured memory mechanisms (Wang et al., 2023a).
+Realistic software tasks may require agents to store and reason over evolving states, feedback logs, intermediate plans, and prior actions (Qian et al., 2025).
 Without hierarchical and queryable memory systems, agents risk repeating errors, forgetting past successes, or producing inconsistent results.
-Emerging solutions such as retrieval-augmented generation and memory summarization offer partial relief, but they remain inadequate for complex, multi-session workflows (maharana2024evaluating, ).
+Emerging solutions such as retrieval-augmented generation and memory summarization offer partial relief, but they remain inadequate for complex, multi-session workflows (Maharana et al., 2024).
 Future research can explore memory architectures that differentiate short-term interactions, mid-term subgoals, and long-term domain knowledge.
 6. Opportunities and Future Directions
+Software Compatibility (Section 6.1) Structured feedback on failures Direct IR interaction Standardized compiler interfaces How do coding agents integrate with tools?
+Context Management (Section 6.2) Structured storage and recall of evolving task information Conditioned retrieval from task state and tool feedback Program state tracing and replay How do coding agents manage memory and contextual information?
+Benchmarks (Section 6.3) Heavily biased training corpus Large absence of real-world tasks Lack of reflecting interactive and iterative programming workflow How to comprehensively evaluate coding agents?
+Human-AI Collaboration (Section 6.4) Frameworks for responsibility allocation Handle uncertainty; Ecosystem integration Coordination in multi-user environments How do coding agents collaborate with human?
+Domain-specific Agents (Section 6.5) Adaptive prompting, fine-tuning, and behavior modulation Reasoning-loop integration of domain tools and diagnostics Runtime domain detection and adaptation How to develop domain-specific coding agents?
+Safety, Alignment and Trust (Section 6.6) Agent behavior grounding Risk detection and reduction Explainable and self-reflective agents How to ensure safer, ethical, and trustworthy coding agents?
+Multi-Agent Collaboration (Section 6.7) Native protocol support for multi-agent systems Session management for dynamic coordination Semantics-aware routing, conflict resolution, and consistency How to design standardized protocols for multi-agent collaboration?
+System Support (Section 6.8) System resource scheduling (CPU, GPU, memory) Reducing overhead in communication with environments Coordinating tasks and resolving conflicts in collaborative codebases How to enhance system-level support for coding agents?
+Figure 8. Summary of potential future research directions of AI agentic programming.
 AI agentic programming represents a fast-evolving research frontier that intersects artificial intelligence, programming languages, and software engineering.
@@
 In this section, we outline several key opportunities and open research directions that can shape the future of this field.
+An overview is provided in Figure 8.
 6.1. Integrating Coding Agents with Tools
@@
 A promising direction is to rethink how programming languages, compilers, and testing frameworks, which are traditionally built for human developers, can be redesigned to support AI coding agents.
-For example, instead of emitting opaque diagnostics, compilers could provide structured feedback explaining why certain optimizations (e.g., vectorization or inlining) fail (maleki2011evaluation, ; ashouri2018survey, ; siso2019evaluating, ).
-These could include semantic barriers like unresolved aliasing, ambiguous data/control flow, or missing annotations (bi2024iterative, ), enabling agents to revise code more precisely.
+For example, instead of emitting opaque diagnostics, compilers could provide structured feedback explaining why certain optimizations (e.g., vectorization or inlining) fail (Maleki et al., 2011a; Ashouri et al., 2018; Siso et al., 2019).
+These could include semantic barriers like unresolved aliasing, ambiguous data/control flow, or missing annotations (Bi et al., 2024), enabling agents to revise code more precisely.
 Beyond diagnostics, compilers can help agents track state across iterations.
@@
 Opening compiler internals also presents a valuable opportunity.
-Coding agents could interact directly with intermediate representations (IRs), such as LLVM IR (lattner2004llvm, ) or MLIR (lattner2021mlir, ), to reason about program structure, verify transformations, or perform static analysis at a semantic level.
+Coding agents could interact directly with intermediate representations (IRs), such as LLVM IR (Lattner and Adve, 2004b) or MLIR (Lattner et al., 2021), to reason about program structure, verify transformations, or perform static analysis at a semantic level.
 Compiler APIs and language servers (e.g., Clang’s LibTooling, the Language Server Protocol) already expose ASTs, symbol tables, and refactoring tools, but a wider adoption may require standardized, introspective interfaces across compilers.
@@
 For instance, agents can dynamically insert instrumentation or launch profiling runs, then use the results to inform optimization choices.
-Coupling these capabilities with autotuning frameworks (chen2018tvm, ; ansel2014opentuner, ) would expand the design space while preserving correctness and safety.
+Coupling these capabilities with autotuning frameworks (Chen et al., 2018a; Ansel et al., 2014) would expand the design space while preserving correctness and safety.
 Finally, advances in structured code representations, such as ASTs, graph-based IRs, and semantic embeddings, offer a foundation for more powerful agent reasoning.
@@
 A key capability of agentic programming lies in managing memory and contextual information across tasks that involve long context reasoning and multiple iterations.
-Unlike traditional code generation, which typically follows a single pass prompt to solution model, agentic workflows for solving real-world software engineering problems involve multiple steps, iterative refinement, and integration with external tools and development environments (cursor, ; wang2024openhands, ; openai2021codex, ; github_copilot, ).
+Unlike traditional code generation, which typically follows a single pass prompt to solution model, agentic workflows for solving real-world software engineering problems involve multiple steps, iterative refinement, and integration with external tools and development environments (cur, 2025; Wang et al., 2024d; OpenAI, 2021; GitHub, 2021).
 Consider an agent tasked with adding a new feature to an open source project, such as implementing a command-line flag to enable verbose logging.
@@
 Agentic programming, therefore, needs mechanisms for memory and context tracking that go beyond simple token limits, enabling coding agents to maintain continuity across extended interactions and tool usage.
-As the memory footprint of LLMs grows linearly with input token length (DBLP:conf/nips/VaswaniSPUJGKP17, ; kaplan2020scaling, ), current LLM-based agents remain constrained by their context windows and lack persistent memory across a long sequence of iterations (schick2023toolformer, ; liu2023lost, ; packer2023memgpt, ).
+As the memory footprint of LLMs grows linearly with input token length (Vaswani et al., 2017; Kaplan et al., 2020), current LLM-based agents remain constrained by their context windows and lack persistent memory across a long sequence of iterations (Schick et al., 2023; Liu et al., 2023b; Packer et al., 2023).
 However, reasoning about real-world programs often requires modeling complex data structures and code context (like function calls) spanning across multiple files, which often exceeds typical context limits.
-Although some industry-scale LLMs claim to support million-token contexts (ding2023longnet, ; team2024gemini, ; ding2024longrope, ), they often rely on random sampling techniques (hosseini2024efficient, ) and fail to leverage program structure or semantics effectively.
+Although some industry-scale LLMs claim to support million-token contexts (Ding et al., 2023; Team et al., 2024; Ding et al., 2024), they often rely on random sampling techniques (Hosseini et al., 2024) and fail to leverage program structure or semantics effectively.
 Therefore, an interesting direction is to design attention mechanisms that are guided by code structure, such as syntax trees, control flow graphs, or data dependencies.
 These structures can help agents focus more accurately on the most relevant parts of a program.
-While approaches like retrieval-augmented generation (RAG) (lewis2020retrieval, ), KV cache offloading (lee2024infinigen, ; sun2024shadowkv, ), and compression (liu2024kivi, ; chang2025palu, ) provide partial solutions, they struggle to provide precise control over long-term dependencies, structured knowledge, and execution histories.
+While approaches like retrieval-augmented generation (RAG) (Lewis et al., 2020), KV cache offloading (Lee et al., 2024; Sun et al., 2024a), and compression (Liu et al., 2024g; Chang et al., 2025) provide partial solutions, they struggle to provide precise control over long-term dependencies, structured knowledge, and execution histories.
 AI coding agents can also benefit from hierarchical memory models that distinguish between short-term interaction history, mid-term planning objectives (such as subgoals and intermediate decisions), and long-term knowledge.
@@
 For example, an agent might summarize a multi-turn session as a sequence of planning decisions and outcomes, highlighting key insights and interventions.
-Another important area is the development of context-aware retrieval strategies that move beyond static similarity-based methods (su2024dragin, ).
+Another important area is the development of context-aware retrieval strategies that move beyond static similarity-based methods (Su et al., 2024).
 During debugging, for instance, an agent could retrieve not only the most recent error message but also similar past failures, proposed fixes, relevant test cases, and their outcomes.
@@
 Table 8. Benchmarks for evaluating LLMs and agentic systems on programming tasks.
-Abbreviations: CP = Competitive Programming, BF = Bug Fixing, FC = Function Completion, CR = CLI Reasoning.
-Benchmark Source Language Task Difficulty Year
-HumanEval Hand-written Python FC Beginner 2021
-MBPP Crowd-sourced Python FC Beginner 2021
-CodeContests CP Python/C++/Java FC Diverse 2022
-HumanEval-X Hand-written Python/C++/Java/JS/Go FC Intermediate 2023
-SWE-Bench GitHub Issues Python BF Expert 2024
-SWE-bench M GitHub Issues JS BF Diverse 2024
-LiveCodeBench CP (live) Python FC, BF Diverse 2024
-TerminalBench Community-curated Shell CR Diverse 2025
-Spider 2.0 Enterprise DB Apps SQL FC Expert 2025
-EffiBench-X Synthetic Python/C++/JS FC, BF Diverse 2025
-Web-Bench Web App Projects JS/TS/HTML/CSS/Python BF, FC Expert 2025
-ProjectEval Open-source repos Python/Java/C++/JS FC, BF, CR Expert 2025
-TRAIL CP Python/Java/C++/JS BF, CR Expert 2025
+Abbreviations: CP = Competitive Programming, BF = Bug Fixing, FC = Function Completion, CR = CLI Reasoning, PO = Performance Optimization.
+[…]
 Note: ”Diverse” difficulty indicates that the benchmark covers a wide range from beginner to expert-level tasks.
-Table 7. AI Intelligence Score (across mathematics, science, coding, and reasoning) and Speed (Output tokens per second) for LLMs on LivecodeBench
-Metric GPT-5 GPT-4o Claude 4 Gemini 2.5 Pro Grok 4 DeepSeek R1 Kimi K2 Qwen3 Solar-Pro Devstral Llama 4 Magistral
-AI score 69 65 59 65 68 59 49 64 43 31 42 36
-Speed 102 148 100 153 76 21 45.2 24 54 44 174 51
 Table 8 summarizes several widely used coding benchmarks for evaluating LLMs and agentic systems on programming tasks.
 These benchmarks vary in their origin, programming language coverage, task types, and difficulty levels.
-Commonly used datasets include HumanEval (openai2021codex, ), HumanEval-X (zheng2023codegeex, ), MBPP (austin2021program, ), SWE-Bench (jimenez2023swe, ), SWE-bench Multimodal (yang2024swe, ), TerminalBench (tbench_2025, ), LiveCodeBench (jain2024livecodebench, ), CodeContests (DBLP:journals/corr/abs-2203-07814, ), Spider2.0 (lei2024spider, ), EffiBench-X (qing2025, ), Web-Bench (xu2025webbench, ), ProjectEval (liu-etal-2025-projecteval, ), and TRAIL (deshpande2025trailtracereasoningagentic, ).
+Commonly used datasets include HumanEval (OpenAI, 2021), HumanEval-X (Zheng et al., 2023b), MBPP (Austin et al., 2021b), SWE-Bench (Jimenez et al., 2023), SWE-bench Multimodal (Yang et al., 2024b), Terminal-Bench (Team, 2025), LiveCodeBench (Jain et al., 2024), CodeContests (Li et al., 2022), Spider2.0 (Lei et al., 2024), EffiBench-X (Qing et al., 2025), Web-Bench (Xu et al., 2025b), ProjectEval (Liu et al., 2025), TRAIL (Deshpande et al., 2025), and GSO (Shetty et al., 2025).
 While these benchmarks have provided valuable insights into the capabilities of LLMs and agentic systems for code generation and bug fixing, they also have important limitations.
-For example, they are heavily biased toward a small set of programming languages - especially Python, which dominates the training corpus of code for current models (twist2025llms, ; chai2024mceval, ).
-This limits the generalizability of evaluation results to domains involving statically typed or domain-specific languages, such as C++ or Rust (twist2025llms, ).
-For example, Table 7 presents the AI Intelligence Score (aggregated across mathematics, science, coding, and reasoning) and the output speed (in tokens per second) for a selection of widely used LLMs on the LiveCodeBench evaluation.
-For the intelligence score, the latest model, GPT-5, achieved 69, indicating that more than 30% of challenges remain unresolved.
+For example, they are heavily biased toward a small set of programming languages - especially Python, which dominates the training corpus of code for current models (Twist et al., 2025; Chai et al., 2024).
+This limits the generalizability of evaluation results to domains involving statically typed or domain-specific languages, such as C++ or Rust (Twist et al., 2025).
+For example, Figure LABEL:fig:res presents the performance of the top-10 LLMs across four representative benchmarks.
+The results reveal clear task-dependent differences.
+For example, all models still perform poorly on software optimization tasks, while they perform well on code generation tasks.
+Overall, performance gaps remain across all benchmarks, indicating that current LLMs still face significant challenges in reasoning, coding, and generalization to complex tasks.
 Furthermore, many existing benchmarks focus on small, self-contained problems that may not be representative of real-world software engineering tasks.
@@
 Another key limitation is that most benchmarks do not capture the interactive, iterative nature of agentic programming.
-In real-world settings, coding agents will need to collaborate with human developers (wang2024openhands, ; cursor, ; openai2021codex, ; github_copilot, ), receive intermediate feedback and confirmation, and rely on external tools such as compilers, debuggers, and test frameworks.
+In real-world settings, coding agents will need to collaborate with human developers (Wang et al., 2024d; cur, 2025; OpenAI, 2021; GitHub, 2021), receive intermediate feedback and confirmation, and rely on external tools such as compilers, debuggers, and test frameworks.
 Benchmarks that assume single-shot or non-interactive task completion fail to reflect the complexity of such multi-step, tool-augmented workflows.
-Addressing these gaps will require the development of more comprehensive and extensible evaluation frameworks.
+Addressing these gaps will require developing more comprehensive and extensible evaluation frameworks.
 Future benchmarks should incorporate realistic tasks that reflect end-to-end development workflows, support multiple programming languages, and enable interaction with tools and human feedback loops.
@@
 6.4. Human-AI Collaboration
-While a long-term vision of agentic programming is to automate the entire software development lifecycle, including writing, debugging, and testing code, near-term opportunities include extending the capabilities of current LLM-based coding assistants (wang2024openhands, ; cursor, ; openai2021codex, ; github_copilot, ).
+While a long-term vision of agentic programming is to automate the entire software development lifecycle, including writing, debugging, and testing code, near-term opportunities include extending the capabilities of current LLM-based coding assistants (Wang et al., 2024d; cur, 2025; OpenAI, 2021; GitHub, 2021).
 Rather than replacing human developers, these systems can act as collaborative partners, supporting workflows in which humans retain strategic oversight.
-Similar to pair programming (hannay2009effectiveness, ), LLM-based agents can assist by proposing ideas, detecting errors, suggesting improvements, and automating routine tasks, thereby augmenting human productivity.
-Designing effective human–AI collaboration models remains an important research challenge.
-This includes developing user interfaces, interaction protocols, and frameworks for allocating responsibilities between human developers and agents, particularly in professional and team-based environments where coordination, trust, and efficiency are essential.
-Unlike traditional code generation tools, agentic systems operate in iterative, feedback-driven loops and are capable of making autonomous decisions based on intermediate results.
-This interaction model shifts the role of the developer from directive command-giver to interactive collaborator, opening up new opportunities and challenges for co-creating software.
-To enable meaningful collaboration, agents must be transparent in their reasoning, responsive to human input, and able to explain their decisions.
-Realizing these capabilities will require advances in interactive prompting (schulhoff2024prompt, ), natural language explanations (creswell2022faithful, ), and context-aware dialogue protocols (shah2025prompt, ; anthropic2024mcp, ).
-One promising research direction is the development of mixed-initiative workflows, in which control shifts fluidly between the developer and the agent.
-In these workflows, the human might specify high-level goals, architectural constraints, or coding conventions, while the agent proactively generates scaffolding code, explores design alternatives, or automates repetitive subtasks.
-The developer can then inspect, accept, reject, or refine the agent’s contributions.
-For instance, an agent could propose multiple refactoring strategies for a large function and generate test cases to verify the correctness of each variant.
-The human collaborator evaluates these options and selects the most appropriate one, potentially with further refinements.
-Supporting such workflows requires robust management of shared context, memory, and goal state across multiple interaction turns.
-Agents must also be able to handle ambiguity and uncertainty.
-Real-world development often involves vague or underspecified requirements.
-Future research may explore agents that can ask clarifying questions (zhang2024modeling, ; mu2024clarifygpt, ), infer intent from surrounding artifacts, or learn user preferences over time.
-For example, when given an instruction like “make this more efficient”, the agent should determine whether to optimize for speed, memory, or readability - potentially by prompting the user for clarification.
-Furthermore, human developers rely on a rich ecosystem of non-code artifacts, such as design diagrams, issue trackers, documentation, and domain-specific knowledge.
-Integrating these contextual signals into the agent’s reasoning pipeline remains an open challenge.
-For instance, when fixing a bug, the agent may need to cross-reference issue tracker discussions or past commits to fully understand the context and constraints of the fix.
-Trust and usability are foundational to collaboration.
-Developers must feel confident that the agent’s actions are safe, interpretable, and reversible.
-Research is needed on techniques that allow agents to explain their decisions, cite relevant documentation or examples (e.g., from internal codebases or public repositories), and highlight trade-offs.
-Users should be empowered to inspect, override, or revert agent-generated changes easily—especially in safety-critical or compliance-sensitive domains such as finance, healthcare, or aerospace.
-Collaboration in team-based environments presents additional opportunities.
-Future systems may support multi-user workflows where agents assist in coordinating tasks across a team, help onboard new developers by answering project-specific questions, or even mediate merge conflicts by suggesting consistent resolutions based on prior history.
-For example, an agent integrated into a version control system might suggest commit messages, flag inconsistent code style across pull requests, or recommend reviewers based on code ownership.
-In summary, enabling rich human-AI collaboration in agentic programming will require progress across multiple dimensions: user interface design, communication strategies, intent modeling, memory management, and trust calibration.
-These systems should be designed not to replace human developers, but to augment their creativity, efficiency, and decision-making.
-Ultimately, effective collaboration will depend on how well agents can adapt to the social and cognitive workflows of modern software development.
+Similar to pair programming (Hannay et al., 2009), LLM-based agents can assist by proposing ideas, detecting errors, suggesting improvements, and automating routine tasks, thereby augmenting human productivity.
+Designing effective models for human-AI collaboration is a key research challenge.
+This involves creating user interfaces, interaction protocols, and responsibility-sharing frameworks suited for professional and team-based environments where coordination, trust, and efficiency are critical.
+Unlike traditional code generation tools, agentic systems operate in iterative, feedback-driven loops and make autonomous decisions based on intermediate results.
+This shifts the developer’s role from command giver to collaborator, opening new opportunities and challenges for co-creating software.
+To support this shift, agents must be transparent in their reasoning, responsive to human input, and capable of explaining their decisions, which calls for advances in interactive prompting (Schulhoff et al., 2024), natural language explanations (Creswell and Shanahan, 2022), and context-aware dialogue protocols (Shah, 2025; Anthropic, 2024a).
+An interesting direction would be mixed initiative workflows, where control moves fluidly between human and agent.
+Developers can specify goals, constraints, or coding conventions, while agents generate scaffolding code, explore alternatives, or automate repetitive tasks.
+The developer then inspects, accepts, rejects, or refines the output.
+For example, an agent may suggest several refactoring options and generate test cases to validate them, leaving the final choice to the human.
+Such workflows must also support interruption and resumption, enabling agents to incorporate new input without losing context.
+This requires toolchains that track shared state across iterations, including goals, partial programs, and transformation history.
+Finally, domain-aware collaboration can strengthen interpretability and safety.
+Agents should handle ambiguous or incomplete requirements by asking clarifying questions (Zhang et al., 2024c; Mu et al., 2024), drawing on design documents, issue trackers, and domain knowledge to refine their reasoning.
+They should also provide clear explanations, cite relevant evidence, and highlight trade-offs to build trust, which is especially critical in domains such as finance, healthcare, and aerospace.
 6.5. Domain Specialization and Adaptability
-General-purpose coding agents often underperform in domain-specific tasks due to a lack of tailored knowledge, tool integration, and contextual understanding (joel2024survey, ; zheng2024well, ).
-While LLMs trained on extensive code corpora can demonstrate general programming capabilities, they may struggle in specialized domains, such as generating Verilog code for hardware synthesis, which involves strict performance constraints, low-level abstractions, or domain-specific libraries and tooling.
-Future research may explore strategies for domain-adaptive prompting, fine-tuning, and agent behavior modulation to improve performance in areas such as embedded systems, data science, scientific computing, high-performance computing, and formal methods.
-For example, in embedded systems programming (englhardt2024exploring, ), agents must reason about memory layout, real-time constraints, and hardware-specific APIs, which are often underrepresented in general training datasets of code.
-Fine-tuning on domain-specific codebases or adapting the agent’s planning strategy to prioritize correctness and safety over code brevity may significantly improve performance.
-Another direction is to build LLM agents that incorporate domain-specific tools and diagnostics into their reasoning loop.
-For instance, a scientific computing agent might integrate with tools like NumPy, MATLAB, or domain-specific simulation engines (openfoam, ; permann2020moose, ; simulink, ), allowing it to generate, run, and analyze numerical experiments autonomously.
-Similarly, an agent for cybersecurity tasks might interact with static analyzers (codeql2020, ), fuzzers (manes2019art, ), or formal verification tools (woodcock2009formal, ) to check code properties as part of its development process.
-Domain specialization can also improve interpretability and safety.
-By narrowing the operational scope of an agent, developers can apply stricter validation, leverage existing formal specifications, and generate more meaningful feedback.
-In safety-critical applications, such as aerospace, automotive, or medical software, agents must operate under rigorous constraints, and tailoring their reasoning and code generation behaviors to domain-specific requirements is essential.
+While LLMs trained on extensive code corpora can demonstrate general programming capabilities (Joel et al., 2024; Zheng et al., 2024a), they may struggle in specialized domains, such as generating Verilog code for hardware synthesis, which involves strict performance constraints, low-level abstractions, or domain-specific libraries and tooling.
+Future research can improve agentic programming by exploring domain adaptive prompting, fine-tuning, and behavior modulation in areas such as embedded systems, data science, scientific computing, high-performance computing, and formal methods.
+In embedded systems programming (Englhardt et al., 2024), for example, agents must reason about memory layout, timing constraints, and hardware-specific APIs, which are often underrepresented in general training data.
+Training on domain-specific code or adjusting planning strategies to prioritise correctness and safety over brevity may enhance performance.
+Integrating domain-specific tools and diagnostics into agents’ reasoning, such as simulation engines (OpenFOAM Foundation, 2023; Permann et al., 2020; The MathWorks, Inc., 2023), static analyzers (Lab, 2020), fuzzers (Manès et al., 2019), or formal verification tools (Woodcock et al., 2009), can further improve robustness.
+Specialisation also supports interpretability and safety, enabling stricter validation, better use of formal specifications, and more reliable feedback in safety-critical domains such as aerospace, automotive, and healthcare.
 Additionally, research into adaptive agent behavior can enable models to detect the domain of a task at runtime and adjust their prompting strategies, tool usage, and explanation styles accordingly.
@@
 6.6. Safety, Alignment, and Trust
-As agentic coding systems become more autonomous and start taking on greater responsibility in software development, it becomes increasingly important to ensure that their behavior aligns with user intent, produces correct results, and avoids unintended changes (ouyang2022training, ; liu2024exploring, ).
-Unlike traditional LLM-based code assistants, agentic systems can take multi-step actions, use external tools, and modify codebases with limited human oversight, making safety and trust essential goals (nvidia2025safetyrecipe, ; hua2024trustagent, ).
+As agentic coding systems start taking on greater responsibility in software development, it becomes increasingly important to ensure that their behavior aligns with user intent, produces correct results, and avoids unintended changes (Ouyang et al., 2022; Liu et al., 2024b).
+Unlike traditional LLM-based code assistants, agentic systems can take multi-step actions, use external tools, and modify codebases with limited human oversight, making safety and trust essential goals (NVIDIA Technical Blog, 2025; Hua et al., 2024).
 One key research direction is building agents that can better understand and follow what users actually want, even when instructions are vague or incomplete.
 Current systems rely heavily on natural language prompts, which can be ambiguous.
-Future work could focus on grounding agent behavior in more structured forms of input, such as constraints, test cases, or high-level goals, that are easier to validate and reason about (fakhoury2024llm, ; huang2023agentcoder, ; li2025structured, ).
-An interesting idea is to design a structured language that developers can use to express their intent clearly.
+Future work could focus on grounding agent behavior in more structured forms of input, such as constraints, test cases, or high-level goals, that are easier to validate and reason about (Fakhoury et al., 2024; Huang et al., 2023; Li et al., 2025b).
+Another interesting idea is to design a structured language that developers can use to express their intent clearly.
 This language would serve as a kind of programming interface for LLMs, allowing users to define what the agent should do, what it must avoid, and what counts as a valid solution.
 For example, a developer might specify that a function’s output must remain the same after refactoring, and the agent would only explore changes that meet this requirement.
-This kind of structure could also make it easier to apply lightweight verification tools, such as static analyzers or type checkers, to ensure the agent’s suggestions are safe and correct (codeql2020, ; li2024iris, ; nunez2024autosafecoder, ).
+This kind of structure could also make it easier to apply lightweight verification tools, such as static analyzers or type checkers, to ensure the agent’s suggestions are safe and correct (Lab, 2020; Li et al., 2024a; Nunez et al., 2024).
 In safety-critical domains like healthcare, finance, or automotive software, agents will also need to follow strict coding standards, legal rules, and certification guidelines.
-Agents could be trained to recognize such constraints or be paired with rule-based systems to flag violations and suggest compliant alternatives (ouyang2022training, ; he2024instruction, ).
+Agents could be trained to recognize such constraints or be paired with rule-based systems to flag violations and suggest compliant alternatives (Ouyang et al., 2022; He et al., 2024a).
 Another important challenge is ethical alignment.
-Because agentic systems are trained on large and diverse codebases, they may learn unsafe or biased practices (liu2023uncovering, ; pearce2025asleep, ; huang2024bias, ).
+Because agentic systems are trained on large and diverse codebases, they may learn unsafe or biased practices (Liu et al., 2023a; Pearce et al., 2025; Huang et al., 2024g).
 Research efforts are needed to detect and reduce risks such as generating insecure code, leaking sensitive data, or reinforcing stereotypes.
 Techniques from responsible AI, like behavioral audits, adversarial testing, and human feedback, can be adapted to this setting.
-To build user trust, agents should also be able to explain their decisions (rajani2019explain, ; creswell2022faithful, ; DBLP:conf/nips/ShinnCGNY23, ).
-Developers need to know why an agent made a change, where the idea came from, and what trade-offs are involved.
-Research on explanation generation, source citation, and visualization of code changes will help make agent behavior more transparent and understandable (liu2024reliability, ; sun2022investigating, ; huang2023citation, ).
-Agents should also be aware of their own limitations.
-When they are uncertain, they should be able to flag their confidence level (kuhn2023semantic, ; creswell2022faithful, ), suggest multiple options (DBLP:conf/nips/MadaanTGHGW0DPY23, ; DBLP:conf/nips/ShinnCGNY23, ), or ask the user for confirmation (mu2024clarifygpt, ; zhang2024modeling, ).
-Designing agents that can adapt their behavior based on task difficulty or user feedback is a promising direction (DBLP:conf/nips/ShinnCGNY23, ; sun2023adaplanner, ; damani2024learning, ; manvi2024adaptive, ; chen2024magicore, ).
-Finally, safety mechanisms like undo and audit trails will be essential.
-Agents should keep track of what they have changed and allow users to roll back actions easily (patil2024goex, ; jiang2024rocode, ).
+To build user trust, agentic programming systems should also be able to explain their decisions (Rajani et al., 2019; Creswell and Shanahan, 2022; Shinn et al., 2023).
+Developers want to know why an agent made a change, where the idea came from, and what trade-offs are involved.
+Research on explanation generation, source citation, and visualization of code changes will help make LLM behavior more transparent and understandable (Liu et al., 2024d; Sun et al., 2022; Huang and Chang, 2023).
+Agents should also be aware of their limitations.
+When they are uncertain, they should be able to flag their confidence level (Kuhn et al., 2023; Creswell and Shanahan, 2022), suggest multiple options (Madaan et al., 2023; Shinn et al., 2023), or ask the user for confirmation (Mu et al., 2024; Zhang et al., 2024c).
+Designing agents that can adapt their behavior based on task difficulty or user feedback is a promising direction (Shinn et al., 2023; Sun et al., 2023b; Damani et al., 2024; Manvi et al., 2024; Chen et al., 2024d).
+Furthermore, safety mechanisms like undo and audit trails will be essential.
+Agents should keep track of what they have changed and allow users to roll back actions easily (Patil et al., 2024; Jiang et al., 2024a).
 Future work could explore automatic snapshotting, reversible code edits, and tight integration with version control systems to support safe collaboration.
+6.7. Multi-agent Collaboration
+As coding agents increasingly operate in teams and interact with heterogeneous services, effective multi-agent collaboration requires standardized communication protocols that provide low-latency, scalable, and semantics-aware coordination.
+However, current solutions remain limited and fall short of these needs.
+Future communication protocols should move beyond ad-hoc adapters and heterogeneous web services toward native support for multi-agent collaboration.
+Promising directions include the design of lightweight, low-latency group communication frameworks that natively support both agent–to–agent and agent–to–tool interactions, while offering standardized session management for dynamic participation and coordination (Xu et al., 2012; A2A, 2025; Verma et al., 2015).
+Recent standards efforts such as IEEE P3394 propose a universal framework with session, channel, and message abstractions, enabling consistent semantics and interoperability across diverse transports (Tong et al., 2025).
+Building on these initiatives, future research should emphasize semantics-aware routing, conflict resolution, and consistency guarantees in distributed multi-agent settings.
+Ultimately, advancing collaboration-centric protocols will be essential for scaling efficient agentic systems in practice.
+6.8. System Support for AI Coding Agents
+Realizing agentic programming at scale also requires strong systems-level support.
+When offered as a cloud service (Amazon, 2025; Microsoft, 2025; KServe Community, 2025; Hugging Face, 2025), LLMs and coding agents must be backed by infrastructures that can efficiently manage resources, communication, and state.
+Agentic workflows combine code generation, compilation, testing, and tool orchestration, with highly variable resource demands (Yang et al., 2024b; Wang et al., 2024d).
+Systems must schedule CPU, GPU, and memory adaptively, balance throughput across users, and minimize latency for interactive tasks (Yu et al., 2021; Fu et al., 2024; Yu et al., 2022; Patel et al., 2024; Aminabadi et al., 2022).
+During the iterative code generation and testing process, agents are likely to exchange large volumes of data with tools like compilers, debuggers, profilers, and repositories.
+Reducing overhead through caching (Kwon et al., 2023; Liu et al., 2024a), compression (Yang et al., 2024a; Jiang et al., 2023; Hooper et al., 2024; Zhao et al., 2024), and lightweight protocols (Google, 2015; Moritz et al., 2018) is essential for efficiency in distributed environments.
+Equally important is context and state management.
+To support interruption, resumption, and collaboration, systems must persist not only source code but also goals, annotations, and intermediate representations.
+Efficient checkpointing and synchronization are key enablers here.
+Finally, multi-agent collaboration introduces distributed systems challenges: coordinating tasks, resolving conflicts, and ensuring consistency when multiple agents - or human–agent teams - work concurrently on shared codebases.
 7. Conclusion
 We have presented a comprehensive review of AI agentic programming.
-This new software paradigm, driven by the success of large language models (LLMs), is a transformative shift in how software can be created, maintained, and evolved.
+This new software paradigm, driven by the recent success of LLMs, is a transformative shift in how software can be created, maintained, and evolved.
 By combining the capabilities of LLMs with planning, tool use, and iterative refinement, coding agents are beginning to automate complex, multi-stage programming workflows that traditionally required significant human involvement.
 These systems can not only generate and test code but also interact with development tools, decompose tasks, and adapt based on feedback, bringing us closer to the vision of autonomous software development.
-In this article, we have introduced a taxonomy of AI coding agents and architectures, reviewed underpinning techniques like context management and tool integration, and discussed how existing benchmarks evaluate the capabilities of coding agents.
+We have introduced a taxonomy of AI coding agents and architectures, reviewed underpinning techniques like context management and tool integration, and discussed how existing benchmarks evaluate the capabilities of coding agents.
 We have summarized the progress made in enabling LLM-based AI agents to reason over tasks, interface with software development tools, and operate in increasingly sophisticated ways.